### v1

# Delayed takedown of illegal content on social media makes moderation ineffective

Abstract
Social media platforms face legal and regulatory demands to swiftly remove illegal content, sometimes under strict takedown deadlines. However, the effects of moderation speed and the impact of takedown deadlines remain underexplored. This study models the relationship between the timeliness of illegal content removal and its prevalence, reach, and exposure on social media. By simulating illegal content diffusion using empirical data from the DSA Transparency Database, we demonstrate that rapid takedown (within hours) significantly reduces illegal content prevalence and exposure, while longer delays decrease the effectiveness of moderation efforts. While these findings support tight takedown deadlines for content removal, such deadlines cannot address the delay in identifying the illegal content and can adversely affect the quality of content moderation.
Introduction
Illegal content, such as copyright violations, child sexual abuse material, and incitement of violence and terrorism, poses serious harm and requires effective moderation to prevent its spread (1, 2). Platforms use various tools for moderation. Interventions like content labels (3, 4), accuracy reminders (5, 6, 7, 8), and designs that add friction to content sharing (9, 10) encourage users to critically engage with content. Other approaches reduce the visibility of harmful content by limiting its searchability, recommendation, and engagement (11, 12). However, for illegal content, laws usually demand strict interventions, such as the removal of illegal content once platforms become aware of it.
Despite existing rules, platforms have often been criticized for failing to act on illegal content, even under government and court orders (13, 14), raising widespread concerns about their commitment to and effectiveness of content moderation (15). In response, various countries have adopted regulations imposing legal requirements on content moderation. Germany led early efforts by passing the Network Enforcement Act (NetzDG) in 2017, followed by a 2020 act to fight hate on the Internet in France (Loi Avia), and Austria’s Communication Platform Law (KoPl-G) and Australia’s Online Safety Act in 2021. In the United States, the proposed Digital Services Oversight and Safety Act (H.R.6796) was rejected in 2023. At the European Union level, the Digital Services Act (DSA) was enacted in 2022, replacing national regulations. All of these recent laws contain specific provisions for content moderation, typically requiring platforms to operate complaint procedures, implement measures against illegal content, and publish transparency reports on the volume of complaints and content moderation decisions (16).
While these regulations share the goal of moderating illegal content, they diverge in their specific requirements for takedown deadlines. For instance, NetzDG and KoPl-G, now replaced by the DSA, mandated immediate content review, requiring platforms to remove “obviously illegal content” within 24 hours and other illegal content within seven days. In contrast, the DSA sets no explicit deadlines, only requiring platforms to make moderation decisions in a “timely manner” (Art. 16 (6)), and to “act expeditiously” to remove or disable access to illegal content (Art. 6 (1)).
From a governance perspective, these regulations raise key questions: How quickly do platforms act in practice, how does moderation speed affect the spread of illegal content, and are takedown deadlines an effective regulatory instrument? These questions remain largely unexplored.
While immediate content removal is intuitively the most effective approach, there is limited research on moderation speed, compliance, and effects. Modeling suggests that moderation delays interfere with intervention effectiveness (17). It was shown, for example, that a four-hour delay reduces prevalence by 55.6% compared to 93.8% for removal within 30 minutes (18). However, rapid removal may reduce precision, causing false positives and concerns about over-moderation or censorship (19, 20, 21). In addition, most existing studies focus on misinformation, such as conspiracy theories, which may not generalize to illegal content, and often examine specific emergencies, such as COVID-19 anti-vaccination narratives (22) or the January 6 Capitol Riot (23), where platforms adopted exceptional measures (24). To inform legal frameworks, it is crucial to also study platform standard practices during routine, non-emergency periods.
Measuring the impact of content-removal requirements at scale is challenging due to limited and inconsistent data. Although platforms self-report moderation speed in community standard enforcement reports, they often use varying metrics for measurement. More standardized data are available under transparency requirements from NetzDG, KoPl-G, and DSA, but scrutiny of these data remains rare (25).
In this paper, we examine how different takedown delays affect the spread of illegal content. Using an agent-based model parameterized with empirically estimated time delays from the DSA Transparency Database (DSA-TDB) (26), we measure the reduction in the prevalence of illegal content and exposure to it as a function of takedown delays. While longer delays predictably result in more users being exposed to illegal content, the exact, non-linear relationship between the two dynamics of takedown and exposure is elucidated by the experiments. The results indicate a steep decline in the effectiveness of moderation as time delay increases. Prompt removal of illegal content significantly reduces visibility compared to systems with no or slow moderation. However, the impact of content removal becomes negligible as the delay extends beyond a couple of weeks. In discussing these findings from a regulatory perspective, we note that while tight content removal deadlines could potentially contribute to preventing the spread of illegal content, they cannot address the time before illegal content is detected. Tight deadlines may also compromise the quality of content moderation.
Results
Modeling illegal content diffusion
We use SimSoM (27), an agent-based model that mimics information diffusion on social networks such as Twitter/X, Mastodon, Bluesky, Threads, or Instagram, where follower relationships need not be reciprocal. Our simulations are conducted on a follower network derived from empirical Twitter data (see Methods). At each time step, a user either posts a new message or reposts a message from their newsfeed. The likelihood that a message is selected by a user for reposting is determined by three factors: the message’s appeal, its social engagement (number of reshares), and its recency. The new or reposted message appears on the newsfeeds of the user’s followers. We use empirical data to model heterogeneous user activity; most agents post rarely, and only a few are hyperactive (see Fig. 1a and Methods).
Figure 1: User activity distribution and probability of posting illegal content. (a) Complementary cumulative distribution of daily average Twitter user activity based on 81,633,118 posts by 37,882,999 accounts during nine days in January 2023 (orange line) and a power-law fit (dashed line). (b) Histograms of illegal content posting probability for low-risk (orange) and high-risk (purple) users, for $p=10^{-4}$ (top), $p=10^{-3}$ (middle), and $p=10^{-2}$ (bottom), based on a large random sample.
We model the legality of a message as a simple binary flag and assume it to be independent from its appeal. This means that legal and illegal messages in the same feed have the same chance of being reshared.
We further assume that illegal content is generated in two ways: (i) by a small fraction $s_{H}$ of “high-risk accounts” (H) that frequently share illegal content, such as child pornography, scams, or terrorist propaganda; and (ii) by the rest of the network, “low-risk accounts” (L) that might occasionally violate the law, such as sharing copyrighted content. We set $s_{H}=0.1$ to reflect a minority of high-risk accounts. We show that the results are robust to this parameter choice in Section Robustness.
We model the probability of posting illegal messages by users in the two groups using beta distributions (see Methods), from which we calculate the illegal content probability $p$. Fig. 1b plots the probability distributions of posting illegal messages by each group for different values of $p$. We use $p=10^{-2}$ and show in Section Robustness that the effectiveness of moderation does not critically depend on this parameter.
To understand the dynamics of illegal content removal, let us consider the lifespan of such content on the platform. While some illegal content is immediately detected and removed using automated detection and decision systems, other is missed due to detection limitations. In such cases, content moderation is delayed until platforms either detect the content themselves later or receive external notification. Additional delays may occur when platforms assess the content against their community standards and national laws. If a violation is confirmed, either the content is removed globally or access is restricted in certain jurisdictions.
We model takedown delay by making the simplifying assumption that at any time step, there is a constant probability $p_{s}$ that a piece of illegal content is not detected and removed, and therefore survives. This yields an exponential survival curve for the delay $t$, $P(t)\sim 2^{-t/\tau}$, where the illegal content half-life $\tau$ is the time taken for half of the illegal content to be removed (median removal time since publication). We use $\tau$ to parameterize content takedown delay. For easier interpretation, we can alternatively define the expected time delay as $\tau/\ln 2$. Since $P(t)\approx p_{s}^{t}$ in the simulation, we set $p_{s}=2^{-1/\tau}$. When a message is removed, it is deleted from all user feeds.
Calibration
We wish to derive a realistic range of $\tau$ values. Effective model calibration depends on reliable empirical data on illegal content takedown delay. We extract Statements of Reason (SoRs) about illegal content removal from the DSA Transparency Database for five major social media platforms: Facebook, Instagram, YouTube, TikTok, and Snapchat (see Methods).
Fig. 2 shows empirical survival time curves for the five social media platforms from DSA-TDB SoR records. By fitting exponential survival time curves $P(t)\sim 2^{-t/\tau}$ to this empirical data, we estimate that the half-life is $\tau\approx 21$ days for Facebook, 60 days for Instagram, 198 days for YouTube, 4 days for TikTok, and 94 days for Snapchat.
The exponential survival curves (Fig. 2 top) fit the data reasonably well for short delays, less for extended times. We can get an intuition for this by noting that the longer the survival of an illegal piece of content, the lower its visibility, and therefore the lower its likelihood to be reported and removed — $p_{s}$ increases instead of remaining constant. To better capture the tails of the data, we also estimate $\tau$ by fitting the linear-log transformed data to $\log_{2}P(t)\sim-t/\tau$. This procedure yields longer half-life estimates $\tau\approx 390$ days for Facebook, 340 days for Instagram, 786 days for YouTube, 1,221 days for TikTok, and 360 days for Snapchat (Fig. 2 bottom). These results highlight the heterogeneity in removal delays within and across platforms.
Figure 2: Distributions of time delay in illegal content takedown on five platforms. Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line). Top: Fit obtained from the raw data. Bottom: Fit obtained from a linear-log transformation of the data to better capture the tails of the curves.
While none of the exponential survival curves perfectly fit the empirical takedown delay data, they provide reasonable bounds for the half-life parameter. Thus, in our experiments, we model various takedown scenarios by exploring a broad range of illegal content half-life values, $2^{-4}\leq\tau\leq 2^{9}$.
Effects of time delay
We measure the effects of time delay in content takedown by the reduction of illegal content throughout the network, which can be quantified by multiple metrics. Prevalence measures the number of posts containing illegal content that are in the system at steady state (see Methods). The more illegal content is present in the system, the more harm it can potentially cause. However, a piece of illegal content may be removed from a user’s feed before the user sees it. Therefore we also measure harm more directly using two exposure metrics. Impressions measures the number of times illegal content is seen by users. Note that a single user can be exposed to many posts with illegal content, or even the same post multiple times. Reach is a complementary measure of exposure, capturing the number of unique users who see any amount of illegal content. For each of these measures, we calculate the reduction percentage relative to a baseline measure without any content removal (see Methods).
Figure 3: Impact of takedown delay on illegal content. The reduction in the prevalence (pink), reach (green), and impressions (purple) of illegal content is plotted as a function of the expected takedown delay ($\tau/\ln 2$). Shading represents the 95% confidence intervals estimated using non-parametric bootstrapping. The vertical lines indicate the expected takedown delays obtaining by fitting the DSA data to exponential curves for different platforms.
Fig. 3 plots the reduction in illegal content prevalence, impressions, and reach as a function of the expected takedown delay, calculated by $\tau/\ln 2$. Content removal has maximal impact under rapid takedowns. For delay below nine hours ($\tau<0.25$ days), removal results in approximately 95–100% reduction of illegal content across all measures.
Moderation becomes less effective as the time delay increases, allowing illegal content to spread before being taken down. Between approximately 9 hours–46 days ($0.25\leq\tau\leq 32$), the reduction in prevalence of illegal content drops from 95% to negligible levels. The half-life values measured from DSA-TDB data for TikTok and Facebook are within this interval. Those for Instagram, Snapchat, and YouTube correspond to even longer delays, where removal has no significant effect on illegal content prevalence. The half-lives obtained by fitting the tails of the platform delay distributions are even longer (see Section Calibration). Although the overall trend is similar across the two exposure metrics, the reduction in impressions and reach drops even faster for delays of up to a few days. For example, when illegal content is removed within a day on average, reach and impressions are only reduced by about 60%. In summary, these findings show that delays beyond a day significantly diminish the effect of content removal.
Robustness
The results presented above may depend on the values of $s_{H}$ (proportion of high-risk accounts), $p$ (illegal content probability), and the underlying follower network. We therefore investigate if the results remain valid when changing these parameters. For these robustness experiments, we use two different values of illegal content half-life, $\tau=2$ and $\tau=8$, and assign users to the high- and low-risk groups randomly in each run.
First, we run experiments to examine the effect of having smaller high-risk groups ($10^{-3}\leq s_{H}\leq 10^{-1}$), which may better reflect realistic scenarios. The main outcome variable, reduction in illegal content prevalence, shows no statistically significant differences (Fig. 4), suggesting that our findings are robust to changes in $s_{H}$.
Figure 4: Robustness of results with respect to different high-risk group sizes $s_{H}$ for $\tau=2$ (top) and $\tau=8$ (bottom) based on 20 simulations. The illegal content reduction does not vary significantly ($P\geq 0.72$ across pairs of experiments with different $s_{H}$ values, using Mann–Whitney U tests with Bonferroni correction).
Table 1: SoRs about illegal content on Facebook, Instagram, YouTube, TikTok, and Snapchat derived from DSA-TDB records between January 1 and December 31, 2024. Note that it is easy for platforms to justify moderation of illegal content with violations of community standards, therefore illegal content may be underreported in the DSA-TDB.
SoRs Illegal SoRs Illegal Ratio
Facebook 487,539,604 8,613 0.00002
Instagram 62,044,832 3,005 0.00005
YouTube 107,390,001 408,400 0.00380
TikTok 977,896,256 93,358 0.00001
Snapchat 4,068,586 2,142 0.00053
Second, we wish to explore how the main result is affected by the volume of illegal content, which varies across platforms and countries. To estimate a plausible range for $p$, let us analyze two sources of empirical data. Table 1 shows that the fraction of illegal SoRs based on the DSA-TDB ranged from 0.00001–0.00380. We can also obtain insights from the Future of Free Speech policy report (28). The authors analyzed comment data collected from Facebook and YouTube across Germany, France, and Sweden during two weeks in June–July, 2023. They report the fraction of comments that disappear ($\frac{\text{disappear}}{\text{total}}$) and the proportion of disappeared comments that are legal ($\frac{\text{legal}}{\text{disappear}}$). From this, we estimate the fraction of comments that are illegal: $\frac{\text{illegal}}{\text{total}}=\frac{\text{disappear}}{\text{total}}\times\frac{\text{illegal}}{\text{disappear}}=\frac{\text{disappear}}{\text{total}}\times(1-\frac{\text{legal}}{\text{disappear}})$. Table 2 shows that this apprach yields a range between 0.00002–0.009.
Table 2: Data from the Future of Free Speech policy report and derived ratios of illegal content.
Disappeared/Total Legal/Disappeared Illegal Ratio
Germany (FB) 0.006 0.997 0.00002
Germany (YT) 0.115 0.989 0.001
France (FB) 0.012 0.921 0.0009
France (YT) 0.072 0.875 0.009
Sweden (FB) 0.005 0.946 0.0003
Sweden (YT) 0.041 0.946 0.002
Based on these estimates, we run experiments with $10^{-4}\leq p\leq 10^{-2}$ (Values $p<10^{-4}$ would require a network with infeasibly large size, as discussed in Methods.) In our main analysis we report the reduction in illegal content prevalence, $1-\frac{\bar{I}_{\text{removal}}}{\langle\bar{I}_{\text{baseline}}\rangle}$. To show that this ratio is independent of $p$, we plot the scaling relationship between $\bar{I}_{\text{removal}}$ and $\bar{I}_{\text{baseline}}$ in Fig. 5a for different $p$ values from Table 2. We observe only a weak sublinear relationship, $\bar{I}_{\text{removal}}\sim\bar{I}_{\text{baseline}}^{0.94}$, suggesting that the ratio is nearly independent of $p$. We also plot in Fig. 5b the reduction in the prevalence of illegal content for different values of $p$ from Table 2. None of the differences between the values in our main results ($p=0.01$) and those for smaller $p$ values are statistically significant. These results suggest that our main results are robust with respect to the volume of illegal content in the system.
Figure 5: Robustness analysis for illegal content probability. (a) Scaling between illegal content prevalence in removal and baseline conditions. Each data point represents the proportion of illegal content in the system with and without removal. Results are based on 40 simulations for each value of $p$ from Table 2, using $\tau=2$. The gray line represents a least-squares fit of the log-transformed data, yielding a slope of $0.94\pm 0.01$. The vertical lines mark the illegal content probabilities from Table 2. (b) Illegal content reduction for different illegal content probabilities corresponding to the same countries and platforms, based on 10 simulations with $\tau=2$ (top) and $\tau=8$ (bottom). The result does not vary significantly ($P\geq 0.13$ across pairs of experiments with different $p$ values, using Mann–Whitney U tests with Bonferroni correction).
Last, we examine the effect of network structure by comparing experiments using the empirical network and a synthetic network (see Methods). There is no statistically significant difference in the reduction in illegal content prevalence between these two networks, indicating the robustness of our findings (Fig. 6).
Figure 6: Robustness analysis with respect to network structures based on 10 simulations with $\tau=2$ (left) and $\tau=8$ (right). The illegal content reduction does not vary significantly ($P=0.11$ and $P=0.22$, respectively, using Mann–Whitney U tests with Bonferroni correction).
Discussion
This paper examines empirical data on illegal content takedown on multiple social media platforms. Our agent-based model calibrated on this data reveals that removal of illegal content within a one-day window significantly reduces its prevalence and exposure, whereas more prolonged delays allow illegal content to permeate user feeds, diminishing the impact of the moderation. Delays beyond 46 days have no impact on the spread of illegal content.
These findings have practical implications for social media moderation and regulation. By adopting more proactive takedown policies, platforms may reduce moderation delay and limit the spread of illegal content. Strengthening the use of and cooperation with trusted flaggers, who identify and report content violating community guidelines or legal standards, can further enhance content moderation. With human flaggers being a limited resource (29), efforts should be directed towards prioritizing the most time-sensitive cases.
For legislation, the importance of prompt and timely moderation appears to support the argument for implementing tight illegal content removal deadlines, as imposed under NetzDG in Germany and KoPl-G in Austria prior to the DSA. However, the practical impact of such deadlines should be carefully assessed in regulatory debates for two reasons. First, regulatory takedown deadlines can only address the timespan between the notification of illegal content and its removal, not the period during which illegal content circulates in the network undetected. As a result, the impact of tight deadlines on the spread of illegal content is inherently limited. In fact, most platforms already report fast reactions after notifications: a median time below 24 hours according to 2024 DSA Transparency Reports and removal of 80% of illegal content within 24 hours according to 2022 national Transparency Reports for Germany and Austria.
Second, tight takedown deadlines may reduce the time necessary for properly assessing the (un)lawfulness of content, increasing the risk of false moderation decisions and lowering the quality of content moderation. Flagged content requires careful verification, including report reviews and appeal resolutions, before removal decisions can be made. Any regulatory framework must ensure effective moderation of illegal content while remaining practical for platforms, and mindful of potential adverse effects. For example, in France, the Loi Avia was struck down by the Constitutional Council (30) due to concerns about restrictions on freedom of expression (31, 32). These considerations suggest that platforms should remove immediately only obviously illegal content. For cases requiring further assessment, platforms should take time for thorough review. Until the assessment is completed, platforms may implement alternative interventions to reduce the circulation of potentially illegal content.
Our model has several limitations. Certain aspects, such as the user activity levels, are calibrated on data from global Twitter users, while the DSA data focuses on European users. Unknown biases could stem from this combination of data sources. The model further presumes uniform spreading dynamics across various types of illegal content. In reality, the appeal of such content may vary widely; scams may attract a broader audience while human trafficking posts may be aimed at specific communities. Due to the lack of data on such content, our model does not distinguish among classes of illegal content or between illegal and legal content. If illegal content were more appealing, our findings would underestimate the adverse effects of time delay on moderation. For modeling content moderation, we make the simplifying assumption that when illegal content is removed, it is deleted from all user feeds. In practice, however, some content can be illegal only in certain jurisdictions and access can be selectively disabled for users in these countries. Our model does not consider such a differentiation.
The DSA-TDB has its own limitations. The database depends heavily on platform voluntary reporting, leading to inconsistencies across platforms. This reflects a fundamental weakness in self-reporting systems, as moderation decisions are often shaped by specific platform terms of service and community guidelines.
Future research could extend our model by accounting for the fact that users often alter their posting patterns in response to moderation decisions (33, 34). The model could also compare the effects of takedown versus alternative types of interventions, as well as explore combinations of different intervention instruments. Finally, future models should consider the trade-offs between the harm caused by persistent illegal content and the unitended effects of moderation tools, e.g., the limitation of free speech caused by false-positive errors.
Methods
Model extensions
We extend SimSoM to model the heterogeneous activity of social media users, where most agents have low activity levels, and only a few are hyperactive. We model this by assigning an activity level $a$ to each user at the beginning of the simulation. We draw this activity from a power law $P(a)\sim a^{\gamma}$, where we obtain the exponent $\gamma=2.85$ by fitting empirical measurements of average daily posts (Fig. 1a). If $a<1$, the user (re)shares a single message with probability $a$ at each time step; else the user (re)shares $\lfloor a\rfloor$ messages.
We wish to model the probability of posting illegal messages by low- and high-risk accounts. Given the lack of empirical data about this probability, we used the posting of content from low-credibility sources as a proxy. Leveraging news source credibility ratings from NewsGuard (35), we categorized a sample of over 250 thousand Twitter users into a low-credibility and a high-credibility group by comparing the average ratings of sources they shared against two thresholds. We then inspected the distributions for posting low-credibility content by these two groups and found that both could be modeled using the parametric family of beta distributions. Based on this proxy, we model the probability $i$ of posting illegal messages by users in each group (low- and high-risk) with a beta distribution $P(i)\sim i^{\alpha-1}(1-i)^{\beta-1}$, where $\alpha>0$ and $\beta>0$ are parameters that express the mean $\frac{\alpha}{\alpha+\beta}$ and skewness of the distribution for that group.
Given the two group distributions, we can express the average illegal content probability across the full population as $p=s_{H}\frac{\alpha_{H}}{\alpha_{H}+\beta_{H}}+(1-s_{H})\frac{\alpha_{L}}{\alpha_{L}+\beta_{L}}$, where $s_{H}$ is the relative size of the high-risk groups, and the $\alpha$ and $\beta$ subscripts indicate the groups.
Since platforms typically do not provide access to original content, there is insufficient scrutiny to determine the legality of content and limited data on the share of illegal content (for an exception see Wagner et al. (36)). Hence, it is difficult to empirically calibrate the illegal content probability $p$ by estimating the percentage of illegal messages in the system. Therefore, we choose different values of $\alpha$ and $\beta$ to explore a broad range of values $10^{-4}\leq p\leq 10^{-2}$.
Content removal
The illegal content prevalence is defined as the average ratio of illegal content in circulation at time $t$: $I_{t}=\frac{1}{N}\sum_{i=1}^{N}f_{i,t}$ where $f_{i,t}$ is the fraction of illegal content in user $i$’s feed at time $t$ and $N$ is the number of agents.
The simulation ends once the system reaches a steady state, signaled by a convergence in the illegal content prevalence. To this end, we calculate an exponential moving average at time $t$ as $\bar{I}_{t}=\rho\bar{I}_{t-1}+(1-\rho)I_{t}$, where $\rho$ is a parameter regulating the importance of older values. We stop the simulation when the difference between two consecutive values of this moving average is smaller than a threshold, $|\bar{I}_{t}-\bar{I}_{t-1}|/\bar{I}_{t-1}<\epsilon$. The parameters $\rho=0.9$ and $\epsilon=0.0001$ used in the reported simulations were tested to ensure that the system’s illegal content prevalence stabilizes at the steady state.
For each set of parameters, we run 70 simulations without any content removal; these are averaged to obtain baseline metrics. We then run 70 simulations with content removal; the reported reduction is averaged across these runs. For example, the reduction in illegal content prevalence is calculated by $(\langle\bar{I}_{\text{baseline}}\rangle-\bar{I}_{\text{removal}})/\langle\bar{I}_{\text{baseline}}\rangle$, where $\bar{I}$ is the illegal content prevalence at the end of a simulation and $\langle\bar{I}_{\text{baseline}}\rangle$ is the baseline average. All simulations start from random conditions.
Data sources
Under the DSA (Art. 15), platforms must publish Transparency Reports detailing “the median time needed for taking the action.” While platforms comply, reporting practices differ significantly, preventing thorough comparison among platforms. Moreover, median turnaround times in transparency reports only cover the time between notification and removal of content, not the time when problematic content circulates in the network before notification. Finally, DSA transparency reports do not indicate moderation delays specific to illegal content, and platforms disallow external access to verify data quality.
More detailed data, suited for our purpose, is available in the newly introduced DSA Transparency Database, which compiles Statements of Reasons (SoRs) explaining single moderation decisions to affected users. These SoRs include the reasons for moderation, as well as information on dates of posting and moderation decisions. We therefore use this data source to estimate the empirical half-life ($\tau$) values. We collect SoRs from the DSA-TDB between January 01 and December 31, 2024. We capture relevant records using the strictest filter, where the platforms explicitly state that the moderation is due to illegal content. [^1] We select SoRs that have DECISION_GROUND_ILLEGAL_CONTENT in the decision_ground field. The database also includes the value DECISION_GROUND_INCOMPATIBLE_CONTENT for the same field. This can cover content that is illegal, which would be indicated in an optional boolean field incompatible_content_illegal. Since platforms do not use this reporting category consistently, we did not include it in our dataset for calibration. We consider four major social media platforms with a sufficient number of relevant SoRs: Facebook, Instagram, YouTube, TikTok, and Snapchat. We exclude LinkedIn and Pinterest due to a lack of SoRs related to illegal content and Twitter/X due to suspicious patterns revealed in its DSA data, as also highlighted in prior work (37). This results in 8,613 SoRs submitted by Facebook, 3,005 by Instagram, 408,400 by YouTube, 93,358 for TikTok, and 2,142 by Snapchat. The time delay in illegal content removal is defined as the difference between the moderation date (application_date field) and the content posting date (content_date field).
Follower network
Simulations are conducted on a follower network derived from empirical Twitter data. The network connects users obtained from a 10% random sample of public tweets (38). The dataset comprises users who posted links to news articles, with likely automated accounts excluded. The original follower network constructed from this dataset has $58{,}296$ nodes and $10{,}632{,}023$ edges. We reduce the size of this network to speed up the simulations in our main experiment (Section Effects of time delay). We employ $k$-core decomposition to identify $N=10{,}006$ nodes that constitute the $k=94$ core. Subsequently, a random selection of edges is removed to match the density of the original network. This procedure produces $E=1{,}809{,}798$ edges, with each node having an average of roughly 180 friends or followers.
For the robustness analyses varying high-risk group size and illegal content probability (Section Robustness), we need a large follower network to run meaningful scenarios using low values of these parameters ($s_{H}=10^{-3}$ and $p=10^{-4}$). Therefore we use the original network, but remove a random set of edges to speed up the simulations. The average in/out-degree is thus reduced to 18 friends per node, resulting in a network with $N=58{,}296$ nodes and $E=1{,}063{,}202$ edges.
To test the robustness of our results with respect to different network structures, we create a third, synthetic network with features ubiquitous in real-world networks, namely, the presence of hubs and clustering (directed triads). To this end, we use a directed variant of the random-walk growth model (39). We configure the parameters of this model to generate a network comparable in size and average degree to the network used in the main experiment. In particular, we initialize the network with 181 fully connected nodes and assume new nodes have fixed out-degree $k_{out}=180$. New nodes are added by first randomly selecting one node to follow. For the remaining $k_{out}-1$ connections, each new node has an equal probability to follow a random friend of the initially chosen node or a randomly selected node in the network. This results in a network with $N=10{,}006$ nodes and $E=1{,}810{,}905$ edges; each node has on average approximately 180 friends/followers.
Author contributions
BT, SK, FM, FS, NJ, SG, and LL conceptualized the study. FM, NJ, SG, and LL acquired funding. BT, SK, and FM developed the methodology. FS and NJ analyzed the regulatory context and data sources. BT, SK, ES, EV, and GN performed formal analysis. BT and FM developed the software. FM provided computing resources. BT conducted the investigation. BT and SK visualized the data. BT, SK, FM, FS and NJ drafted the manuscript. All authors reviewed and edited the draft.
Data availability
Code and data to implement the model and reproduce results are available at github.com/osome-iu/simsom_removal.
Acknowledgements
We are grateful to Alessandro Flammini, Taekho You, Samuel Groesch, and Azza Bouleimen for helpful discussions; and to Nick Liu for the data collection used to estimate user activities. This work was supported in part by the Swiss National Science Foundation (Sinergia grant CRSII5_209250) and by the Knight Foundation.

### v2

@@
 Abstract
-Social media platforms face legal and regulatory demands to swiftly remove illegal content, sometimes under strict takedown deadlines.
-However, the effects of moderation speed and the impact of takedown deadlines remain underexplored.
-This study models the relationship between the timeliness of illegal content removal and its prevalence, reach, and exposure on social media.
-By simulating illegal content diffusion using empirical data from the DSA Transparency Database, we demonstrate that rapid takedown (within hours) significantly reduces illegal content prevalence and exposure, while longer delays decrease the effectiveness of moderation efforts.
-While these findings support tight takedown deadlines for content removal, such deadlines cannot address the delay in identifying the illegal content and can adversely affect the quality of content moderation.
+Illegal content on social media poses significant societal harm and necessitates timely removal.
+However, the impact of the speed of content removal on prevalence, reach, and exposure to illegal content remains underexplored.
+This study examines the relationship with a systematic analysis of takedown delays using data from the EU Digital Services Act Transparency Database, covering five major platforms over a one-year period.
+We find substantial variation in takedown delay, with some content remaining online for weeks or even months.
+To evaluate how these delays affect the prevalence and reach of illegal content and exposure to it, we develop an agent-based model and calibrate it to empirical data.
+We simulate illegal content diffusion, revealing that rapid takedown (within hours) significantly reduces prevalence, reach, and exposure to illegal content, while longer delays fail to reduce its spread.
+Though the effect of delay may seem intuitive, our simulations quantify exactly how takedown speed shapes the spread of illegal content.
+Building on these results, we point to the benefits of faster content removal to effectively curb the spread of illegal content, while also considering the limitations of strict enforcement policies.
 Introduction
-Illegal content, such as copyright violations, child sexual abuse material, and incitement of violence and terrorism, poses serious harm and requires effective moderation to prevent its spread (1, 2).
-Platforms use various tools for moderation.
-Interventions like content labels (3, 4), accuracy reminders (5, 6, 7, 8), and designs that add friction to content sharing (9, 10) encourage users to critically engage with content.
-Other approaches reduce the visibility of harmful content by limiting its searchability, recommendation, and engagement (11, 12).
-However, for illegal content, laws usually demand strict interventions, such as the removal of illegal content once platforms become aware of it.
-Despite existing rules, platforms have often been criticized for failing to act on illegal content, even under government and court orders (13, 14), raising widespread concerns about their commitment to and effectiveness of content moderation (15).
-In response, various countries have adopted regulations imposing legal requirements on content moderation.
-Germany led early efforts by passing the Network Enforcement Act (NetzDG) in 2017, followed by a 2020 act to fight hate on the Internet in France (Loi Avia), and Austria’s Communication Platform Law (KoPl-G) and Australia’s Online Safety Act in 2021.
-In the United States, the proposed Digital Services Oversight and Safety Act (H.R.6796) was rejected in 2023.
-At the European Union level, the Digital Services Act (DSA) was enacted in 2022, replacing national regulations.
-All of these recent laws contain specific provisions for content moderation, typically requiring platforms to operate complaint procedures, implement measures against illegal content, and publish transparency reports on the volume of complaints and content moderation decisions (16).
-While these regulations share the goal of moderating illegal content, they diverge in their specific requirements for takedown deadlines.
-For instance, NetzDG and KoPl-G, now replaced by the DSA, mandated immediate content review, requiring platforms to remove “obviously illegal content” within 24 hours and other illegal content within seven days.
-In contrast, the DSA sets no explicit deadlines, only requiring platforms to make moderation decisions in a “timely manner” (Art.
-16 (6)), and to “act expeditiously” to remove or disable access to illegal content (Art.
-6 (1)).
-From a governance perspective, these regulations raise key questions: How quickly do platforms act in practice, how does moderation speed affect the spread of illegal content, and are takedown deadlines an effective regulatory instrument?
-These questions remain largely unexplored.
-While immediate content removal is intuitively the most effective approach, there is limited research on moderation speed, compliance, and effects.
-Modeling suggests that moderation delays interfere with intervention effectiveness (17).
-It was shown, for example, that a four-hour delay reduces prevalence by 55.6% compared to 93.8% for removal within 30 minutes (18).
-However, rapid removal may reduce precision, causing false positives and concerns about over-moderation or censorship (19, 20, 21).
-In addition, most existing studies focus on misinformation, such as conspiracy theories, which may not generalize to illegal content, and often examine specific emergencies, such as COVID-19 anti-vaccination narratives (22) or the January 6 Capitol Riot (23), where platforms adopted exceptional measures (24).
-To inform legal frameworks, it is crucial to also study platform standard practices during routine, non-emergency periods.
-Measuring the impact of content-removal requirements at scale is challenging due to limited and inconsistent data.
-Although platforms self-report moderation speed in community standard enforcement reports, they often use varying metrics for measurement.
-More standardized data are available under transparency requirements from NetzDG, KoPl-G, and DSA, but scrutiny of these data remains rare (25).
-In this paper, we examine how different takedown delays affect the spread of illegal content.
-Using an agent-based model parameterized with empirically estimated time delays from the DSA Transparency Database (DSA-TDB) (26), we measure the reduction in the prevalence of illegal content and exposure to it as a function of takedown delays.
-While longer delays predictably result in more users being exposed to illegal content, the exact, non-linear relationship between the two dynamics of takedown and exposure is elucidated by the experiments.
-The results indicate a steep decline in the effectiveness of moderation as time delay increases.
-Prompt removal of illegal content significantly reduces visibility compared to systems with no or slow moderation.
-However, the impact of content removal becomes negligible as the delay extends beyond a couple of weeks.
-In discussing these findings from a regulatory perspective, we note that while tight content removal deadlines could potentially contribute to preventing the spread of illegal content, they cannot address the time before illegal content is detected.
-Tight deadlines may also compromise the quality of content moderation.
-Results
-Modeling illegal content diffusion
-We use SimSoM (27), an agent-based model that mimics information diffusion on social networks such as Twitter/X, Mastodon, Bluesky, Threads, or Instagram, where follower relationships need not be reciprocal.
-Our simulations are conducted on a follower network derived from empirical Twitter data (see Methods).
-At each time step, a user either posts a new message or reposts a message from their newsfeed.
-The likelihood that a message is selected by a user for reposting is determined by three factors: the message’s appeal, its social engagement (number of reshares), and its recency.
-The new or reposted message appears on the newsfeeds of the user’s followers.
-We use empirical data to model heterogeneous user activity; most agents post rarely, and only a few are hyperactive (see Fig.
-1a and Methods).
+Illegal and harmful content — including copyright violations, child sexual abuse material (CSAM), and incitement to violence or terrorism — poses serious harm and demands effective moderation efforts (1, 2).
+Platforms have developed a broad range of moderation tools and strategies that vary in severity.
+For some problematic content, such as misinformation and hate speech, interventions have been designed to promote critical user engagement, such as sensitive content labels (3, 4), accuracy reminders (5, 6, 7, 8), and friction mechanisms in content sharing (9, 10, 11).
+These “soft” interventions alter the environment to nudge user behaviors and are thus contingent on user motivations.
+Other techniques reduce the visibility of problematic content by limiting its searchability, recommendation, and engagement (12, 13, 14).
+Stricter interventions include content removal (takedown), account suspensions, and permanent bans (15, 16, 17).
+These approaches enable swift disruption of content virality and minimize exposure, but also raise concerns about censorship (18).
+Given the trade-offs of different moderation approaches, previous modeling work has shown that combining interventions is an effective approach to leverage the strengths of each while mitigating their limitations (19).
+Nevertheless, for clear-cut cases of illegal content, removal or deactivation in concerned jurisdictions remains a legal demand for social media platforms.
+The importance of timely content removal has also been recognized in regulatory frameworks.
+Laws such as Germany’s Network Enforcement Act (NetzDG), Austria’s Communication Platform Act (KoPl-G), and France’s Loi Avia required platforms to remove “obviously illegal” content within 24 hours and other illegal content within seven days.
+These provisions reflected a clear regulatory emphasis on speed and have drawn criticism both for not being strict enough and for being too strict — potentially infringing on free expression (20, 21, 22, 23).
+One of the critical issues is the trade-off between moderation speed and accuracy: removing content very quickly can reduce its spread but may also lower precision, causing false-positive errors and legitimate concerns about over-moderation.
+The Digital Services Act (DSA) replaced precise deadlines with broader requirements: platforms must act “in a timely manner” (Art.
+16(6)) and remove or disable access to illegal content “expeditiously” (Art.
+6(1)).
+While such flexibility avoids rigid constraints, it also marks a shift away from legally enforceable time limits.
+As legal requirements have shifted, the question of how takedown speed affects the prevalence, reach, and exposure of illegal content becomes highly relevant.
+However, thus far, policies enforcing takedown speed lack empirical grounding, making it difficult for policymakers to design and evaluate appropriate policies.
+This paper provides a comprehensive evaluation of the effects of moderation speed by proposing a model that measures how prevalence, exposure, and reach of illegal content change in response to takedown delay.
+In addition, we scrutinize available data on takedown delay of selected platforms and discuss the consequences and limitations of regulatory takedown deadlines.
+Available research on the speed of content removal shows that delay in response reduces prevalence and exposure (24, 25).
+However, the existing studies largely focus on mis- and disinformation, such as misleading claims or conspiracy theories, which may not generalize to illegal content.
+Moreover, much of this work centers on emergencies, such as COVID-19 anti-vaccination narratives (26) or the January 6 Capitol Riot (25), during which platforms adopted exceptional measures.
+To inform legal frameworks, it is necessary to examine the standard operating procedures of platforms during routine, non-emergency periods.
+Studying the impact of content removal policies at scale is challenging, both because it would be ethically impermissible to experiment with harmful or illegal content, and because data on removals is often inaccessible due to privacy concerns.
+Currently, no studies systematically examine illegal content removal practices across platforms or quantify the effects of takedown delays.
+Based on a year of data from the DSA Transparency Database (DSA-TDB) (27), we show that illegal content removal speed varies widely across platforms, ranging from several days to weeks, and in some cases, up to a year.
+By calibrating an agent-based model with empirically estimated takedown delays, we quantify the relationship between moderation speed and the reduction of illegal content spread.
+While longer delays predictably result in greater prevalence, reach and exposure of illegal content, our simulations reveal the precise, non-linear effect of takedown speed.
+The results indicate a steep increase in prevalence and exposure as takedown delay increases.
+Prompt removal of illegal content markedly reduces spread compared to systems with no or slow moderation.
+The impact of content removal becomes negligible as the delay extends beyond a couple of weeks — a timeframe that, according to our data, is typical for the takedown of illegal content on most platforms.
+Our findings indicate that faster content removal can significantly reduce the spread of illegal material.
+However, current takedown delays often allow such content to circulate widely before action is taken.
+This suggests that regulatory frameworks like the DSA may benefit from stronger enforcement provisions aimed at promoting more timely and effective moderation.
+However, in the Discussion, we point to limitations of takedown deadlines and argue that any legal framework should balance the goal of timely removal with the need to ensure the quality of moderation and account for platform-level operational constraints.
+Methods
+Modeling social media diffusion
+We use SimSoM (28), an agent-based model that mimics information diffusion on social networks such as Twitter/X, Mastodon, Bluesky, Threads, or Instagram.
+The platform is represented as a directed network, where nodes correspond to user accounts and edges denote follower relationships, which need not be reciprocal.
+As in real-world platforms, information circulates through messages that appear on user news feeds.
+At each time step, a user can post new original messages, reshare content from their feeds, or take no action.
+Messages represent information that could take the form of text, links, hashtags, images, or other media.
+To select a message for resharing, an agent samples from the news feed, which contains distinct messages shared by their friends.
+Once shared by an agent, a message appears in the news feeds of their followers, allowing for propagation across the network.
+Thus, the exposure of the message depends on its author’s follower count — agents with more followers will achieve greater exposure.
+Capturing mechanisms of real-world ranking algorithms, agents prefer to reshare messages that are appealing, recent, and popular (29).
+The appeal of a message is a scalar value that captures the intrinsic attractiveness of messages.
+This could derive from properties such as novelty or emotional appeal.
+The distribution of message appeal captures a broad distribution of virality often seen in real-world information diffusion networks (28).
+The recency of a message decreases as time passes since it was first introduced to the agent’s news feed.
+The popularity of a message is the number times it was reshared by the agent’s friends.
+Building on the base model, we incorporate additional features to examine the diffusion of illegal content.
+Specifically, each message includes a binary legality attribute indicating whether the content is legal or illegal.
+Different types of illegal content may be more or less appealing.
+For example, CSAM is criminally and morally condemned, therefore resharing is expected to be rare.
+More common forms of illegal content, such as copyrighted media, tend to be highly engaging.
+Lacking data on the correlation between legality and appeal, we make the simplifying assumption that the two are independent.
 Figure 1: User activity distribution and probability of posting illegal content.
-(a) Complementary cumulative distribution of daily average Twitter user activity based on 81,633,118 posts by 37,882,999 accounts during nine days in January 2023 (orange line) and a power-law fit (dashed line).
-(b) Histograms of illegal content posting probability for low-risk (orange) and high-risk (purple) users, for $p=10^{-4}$ (top), $p=10^{-3}$ (middle), and $p=10^{-2}$ (bottom), based on a large random sample.
-We model the legality of a message as a simple binary flag and assume it to be independent from its appeal.
-This means that legal and illegal messages in the same feed have the same chance of being reshared.
-We further assume that illegal content is generated in two ways: (i) by a small fraction $s_{H}$ of “high-risk accounts” (H) that frequently share illegal content, such as child pornography, scams, or terrorist propaganda; and (ii) by the rest of the network, “low-risk accounts” (L) that might occasionally violate the law, such as sharing copyrighted content.
-We set $s_{H}=0.1$ to reflect a minority of high-risk accounts.
-We show that the results are robust to this parameter choice in Section Robustness.
-We model the probability of posting illegal messages by users in the two groups using beta distributions (see Methods), from which we calculate the illegal content probability $p$.
+(a) Complementary cumulative distribution (CCDF) of daily average Twitter user activity based on 81,633,118 posts by 37,882,999 accounts during nine days in January 2023 (orange line).
+The dashed line provides a scaling reference, with the slope 2.85-2.85 representing the exponent of the power law fit.
+(b) Illegal content posting distributions for low-risk (orange) and high-risk (purple) users, for mean illegal content probability $p=10^{-4}$ (top), $p=10^{-3}$ (middle), and $p=10^{-2}$ (bottom), based on a large random sample.
+We also extend SimSoM to model the heterogeneous activity of social media users, where most agents have low activity levels, and only a few are hyperactive.
+We model this by assigning an activity level $a$ to each user at the beginning of the simulation.
+We draw this activity from a power law $P(a)\sim a^{\gamma}$, where we obtain the exponent $\gamma=2.85$ by fitting empirical measurements of average daily posts (Fig.
+1a).
+If $a<1$, the user (re)shares a single message with probability $a$ at each time step; else the user (re)shares $\lfloor a\rfloor$ messages.
+In the following sections, we outline our approach to examine the relationship between takedown delay and effectiveness.
+We first describe the modeling of illegal content diffusion and takedown, followed by an analysis of the DSA-TDB to estimate expected takedown times, and finally, we detail the experimental setup.
+Modeling illegal content diffusion and removal
+We assume that illegal content can be generated by any user, though only a small fraction $s_{H}$ of “high-risk accounts” (H) frequently share such content — for example, CSAM, scams, or terrorist propaganda.
+The rest of the network, “low-risk accounts” (L) may occasionally violate the law, such as sharing copyrighted content.
+The illegal content probability $p$ represents the aggregated probability across user groups, capturing the average likelihood that a user in the population posts illegal content.
+Specifically, the probability of posting illegal messages by both low- and high-risk accounts follows a beta distribution $P(i)\sim i^{\alpha-1}(1-i)^{\beta-1}$, where $\alpha>0$ and $\beta>0$ are parameters that express the mean $\frac{\alpha}{\alpha+\beta}$ and the skewness of the distribution for that group.
+Given the two group distributions, we can express the average illegal content probability across the full population as $p=s_{H}\frac{\alpha_{H}}{\alpha_{H}+\beta_{H}}+(1-s_{H})\frac{\alpha_{L}}{\alpha_{L}+\beta_{L}}$, where $s_{H}$ is the relative size of the high-risk groups, and the $\alpha$ and $\beta$ subscripts indicate the groups.
 Fig.
 1b plots the probability distributions of posting illegal messages by each group for different values of $p$.
-We use $p=10^{-2}$ and show in Section Robustness that the effectiveness of moderation does not critically depend on this parameter.
-To understand the dynamics of illegal content removal, let us consider the lifespan of such content on the platform.
+Since platforms typically do not provide access to original content, there is insufficient scrutiny to determine the legality of content and limited data on the share of illegal content (for an exception see Wagner et al.
+(30)).
+Hence, it is difficult to empirically calibrate the illegal content probability $p$ by estimating the percentage of illegal messages in the system.
+In our experiments, unless stated otherwise, we set $s_{H}=0.1$ to reflect a minority of high-risk accounts.
+We use the values of $\alpha_{L}=0.1$, $\beta_{L}=90$ for the low-risk group and $\alpha_{H}=3$, $\beta_{H}=30$ for the high-risk group, resulting in $p=10^{-2}$.
+We show that the results are robust to these parameter choices in the Robustness section.
+To understand the dynamics of illegal content removal, let us consider the lifespan of such content on a platform.
 While some illegal content is immediately detected and removed using automated detection and decision systems, other is missed due to detection limitations.
@@
 When a message is removed, it is deleted from all user feeds.
-Calibration
-We wish to derive a realistic range of $\tau$ values.
-Effective model calibration depends on reliable empirical data on illegal content takedown delay.
-We extract Statements of Reason (SoRs) about illegal content removal from the DSA Transparency Database for five major social media platforms: Facebook, Instagram, YouTube, TikTok, and Snapchat (see Methods).
+In the next section we derive a realistic range of $\tau$ values from empirical data for five major social media platforms.
+DSA data analysis
+Figure 2: Distributions (CCDF) of time delay in illegal content takedown on five platforms.
+Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line).
+Top panels: Fits based on raw data with linear y-axes.
+Bottom panels: Fits based on a linear-log transformation of the data, highlighting extreme time delay values.
+Under the DSA (Art.
+15), platforms must publish Transparency Reports detailing “the median time needed for taking the action.” While platforms comply, reporting practices differ significantly, preventing thorough comparison among platforms.
+Moreover, median turnaround times in transparency reports only cover the time between notification and removal of content, not the time when problematic content circulates in the network before notification.
+Finally, DSA transparency reports do not indicate moderation delays specific to illegal content, and platforms disallow external access to verify data quality.
+More detailed data, suited for our purpose, is available in the DSA Transparency Database, which compiles Statements of Reasons (SoRs) explaining single moderation decisions to affected users.
+These SoRs include the reasons for moderation, as well as information on dates of posting and moderation decisions.
+We therefore use this data source to estimate the empirical half-life ($\tau$) values.
+We collect SoRs from the DSA-TDB between January 1 and December 31, 2024.
+We capture relevant records using the strictest filter, where the platforms explicitly state that the moderation is due to illegal content.
+[^1] We select SoRs that have DECISION_GROUND_ILLEGAL_CONTENT in the decision_ground field.
+The database also includes the value DECISION_GROUND_INCOMPATIBLE_CONTENT for the same field.
+This can cover content that is illegal, which would be indicated in an optional boolean field incompatible_content_illegal.
+Since platforms do not use this reporting category consistently, we did not include it in our dataset for calibration.
+We consider five major social media platforms with a sufficient number of relevant SoRs: Facebook, Instagram, YouTube, TikTok, and Snapchat.
+We exclude LinkedIn and Pinterest due to a lack of SoRs related to illegal content and Twitter/X due to suspicious patterns revealed in its DSA data, as also highlighted in prior work (31).
+This results in 8,613 SoRs submitted by Facebook, 3,005 by Instagram, 408,400 by YouTube, 90,409 by TikTok, [^2] We excluded 2,949 observations with invalid input for the content creation date (2000-01-01).
+and 2,142 by Snapchat.
+The takedown delay is defined as the difference between the moderation date (application_date field) and the content posting date (content_date field).
 Fig.
@@
 To better capture the tails of the data, we also estimate $\tau$ by fitting the linear-log transformed data to $\log_{2}P(t)\sim-t/\tau$.
-This procedure yields longer half-life estimates $\tau\approx 390$ days for Facebook, 340 days for Instagram, 786 days for YouTube, 1,221 days for TikTok, and 360 days for Snapchat (Fig.
+This procedure yields longer half-life estimates $\tau\approx 390$ days for Facebook, 340 days for Instagram, 786 days for YouTube, 249 days for TikTok, and 360 days for Snapchat (Fig.
 2 bottom).
-These results highlight the heterogeneity in removal delays within and across platforms.
-Figure 2: Distributions of time delay in illegal content takedown on five platforms.
-Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line).
-Top: Fit obtained from the raw data.
-Bottom: Fit obtained from a linear-log transformation of the data to better capture the tails of the curves.
-While none of the exponential survival curves perfectly fit the empirical takedown delay data, they provide reasonable bounds for the half-life parameter.
-Thus, in our experiments, we model various takedown scenarios by exploring a broad range of illegal content half-life values, $2^{-4}\leq\tau\leq 2^{9}$.
-Effects of time delay
+While none of the exponential survival curves perfectly fit the empirical takedown delay data, the short and long estimates from these different fits provide reasonable bounds for the half-life parameter.
+We therefore use them to determine a range of average delay for each platform.
+Simulation framework and parameters
+Based on the above analysis, we simulate various takedown scenarios by exploring a broad range of illegal content half-life values, $2^{-4}\leq\tau\leq 2^{9}$ days.
+Illegal content prevalence is defined as the average ratio of illegal content in circulation at time $t$: $I_{t}=\frac{1}{N}\sum_{i=1}^{N}f_{i,t}$ where $f_{i,t}$ is the fraction of illegal content in user $i$’s feed at time $t$ and $N$ is the number of agents.
+The simulation ends once the system reaches a steady state, signaled by a convergence in the illegal content prevalence.
+To this end, we calculate an exponential moving average at time $t$ as $\bar{I}_{t}=\rho\bar{I}_{t-1}+(1-\rho)I_{t}$, where $\rho$ is a parameter regulating the importance of older values.
+We stop the simulation when the difference between two consecutive values of this moving average is smaller than a threshold, $|\bar{I}_{t}-\bar{I}_{t-1}|/\bar{I}_{t-1}<\epsilon$.
+The parameters $\rho=0.9$ and $\epsilon=0.0001$ used in the reported simulations were tested to ensure that the system’s illegal content prevalence stabilizes at the steady state.
+For each set of parameters, we run 70 simulations without any content removal; these are averaged to obtain baseline metrics.
+We then run 70 simulations with content removal; illegal content prevalence and other quantities are averaged across these runs.
+All simulations start from random conditions.
+Simulations are conducted on a follower network derived from empirical Twitter data.
+The network connects users obtained from a 10% random sample of public tweets (32).
+The dataset comprises users who posted links to news articles, with likely automated accounts excluded.
+The original follower network constructed from this dataset has 58,29658{,}296 nodes and 10,632,02310{,}632{,}023 edges.
+We reduce the size of this network to speed up the simulations in our main experiment (Effects of takedown delay section).
+We employ $k$-core decomposition to identify N=10,006N=10{,}006 nodes that constitute the $k=94$ core.
+Subsequently, a random selection of edges is removed to match the density of the original network.
+This procedure produces E=1,809,798E=1{,}809{,}798 edges, with each node having an average of roughly 180 friends or followers.
+For the robustness analyses, we vary high-risk group size and illegal content probability (Robustness section).
+We therefore need a large follower network to run meaningful scenarios using low values of these parameters ($s_{H}=10^{-3}$ and $p=3\times 10^{-4}$).
+Therefore, we use the original network but remove a random set of edges to speed up the simulations.
+The average in/out-degree is thus reduced to 18 friends per node, resulting in a network with N=58,296N=58{,}296 nodes and E=1,063,202E=1{,}063{,}202 edges.
+To test the robustness of our results with respect to different network structures, we create a third, synthetic network with features ubiquitous in real-world networks, namely, the presence of hubs and clustering (directed triads).
+To this end, we use a directed variant of the random-walk growth model (33).
+We configure the parameters of this model to generate a network comparable in size and average degree to the network used in the main experiment.
+In particular, we initialize the network with 181 fully connected nodes and assume new nodes have fixed out-degree $k_{out}=180$.
+New nodes are added by first randomly selecting one node to follow.
+For the remaining $k_{out}-1$ connections, each new node has an equal probability to follow a random friend of the initially chosen node or a randomly selected node in the network.
+This results in a network with N=10,006N=10{,}006 nodes and E=1,810,905E=1{,}810{,}905 edges.
+Results
+Takedown delay
+Our analysis of the DSA-TDB for five major social media platforms reveals substantial variation in illegal content takedown delays — ranging from several days to weeks, and in some cases up to a year or more.
+TikTok has the fastest takedown speed, with most illegal content lasting approximately 6–359 days.
+YouTube has the the slowest content takedown time, where illegal content can take between 286 days and 3 years to be removed.
+The average illegal content life time for other platforms falls between these extremes, also ranging widely: approximately 30–563 days for Facebook, 87–491 days for Instagram, and 136–519 days for Snapchat.
+(See DSA data analysis for how these ranges are obtained.)
+The observed heterogeneity in takedown delays may also reflect differences in the types of illegal content being moderated.
+Some content types, such as copyright violations, can often be verified quickly, while others may require more time for investigation.
+Effects of takedown delay
 We measure the effects of time delay in content takedown by the reduction of illegal content throughout the network, which can be quantified by multiple metrics.
@@
 Reach is a complementary measure of exposure, capturing the number of unique users who see any amount of illegal content.
-For each of these measures, we calculate the reduction percentage relative to a baseline measure without any content removal (see Methods).
+For each of these measures, we calculate the reduction percentage relative to a baseline measure without any content removal.
+For example, the reduction in illegal content prevalence is calculated by $(\langle\bar{I}_{\text{baseline}}\rangle-\bar{I}_{\text{removal}})/\langle\bar{I}_{\text{baseline}}\rangle$, where $\bar{I}$ is the illegal content prevalence at the end of a simulation and $\langle\bar{I}_{\text{baseline}}\rangle$ is the baseline average.
 Figure 3: Impact of takedown delay on illegal content.
@@
 Shading represents the 95% confidence intervals estimated using non-parametric bootstrapping.
-The vertical lines indicate the expected takedown delays obtaining by fitting the DSA data to exponential curves for different platforms.
+The vertical lines indicate the expected takedown delays obtained by fitting the DSA data to exponential curves for different platforms.
 Fig.
-3 plots the reduction in illegal content prevalence, impressions, and reach as a function of the expected takedown delay, calculated by $\tau/\ln 2$.
+3 plots the reduction in illegal content prevalence, impressions, and reach as a function of the expected takedown delay.
 Content removal has maximal impact under rapid takedowns.
-For delay below nine hours ($\tau<0.25$ days), removal results in approximately 95–100% reduction of illegal content across all measures.
-Moderation becomes less effective as the time delay increases, allowing illegal content to spread before being taken down.
-Between approximately 9 hours–46 days ($0.25\leq\tau\leq 32$), the reduction in prevalence of illegal content drops from 95% to negligible levels.
-The half-life values measured from DSA-TDB data for TikTok and Facebook are within this interval.
-Those for Instagram, Snapchat, and YouTube correspond to even longer delays, where removal has no significant effect on illegal content prevalence.
-The half-lives obtained by fitting the tails of the platform delay distributions are even longer (see Section Calibration).
+For expected delay below nine hours (corresponding to half-life $\tau<0.25$ days), removal results in approximately 95–100% reduction of illegal content across all measures.
+Moderation becomes less effective as the time delay increases, allowing illegal content to spread before being removed.
+Between approximately 9 hours–46 days ($0.25\leq\tau\leq 32$ days), the reduction in prevalence of illegal content drops from 95% to negligible levels.
+The empirical values measured from DSA-TDB data for TikTok and Facebook are within this interval.
+Those for Instagram, Snapchat, and YouTube correspond to longer delays, where removal has no significant effect on illegal content prevalence.
+The high ends of the delay ranges are even longer, as mentioned in the previous subsection.
 Although the overall trend is similar across the two exposure metrics, the reduction in impressions and reach drops even faster for delays of up to a few days.
@@
 The illegal content reduction does not vary significantly ($P\geq 0.72$ across pairs of experiments with different $s_{H}$ values, using Mann–Whitney U tests with Bonferroni correction).
-Table 1: SoRs about illegal content on Facebook, Instagram, YouTube, TikTok, and Snapchat derived from DSA-TDB records between January 1 and December 31, 2024.
-Note that it is easy for platforms to justify moderation of illegal content with violations of community standards, therefore illegal content may be underreported in the DSA-TDB.
-SoRs Illegal SoRs Illegal Ratio
-Facebook 487,539,604 8,613 0.00002
-Instagram 62,044,832 3,005 0.00005
-YouTube 107,390,001 408,400 0.00380
-TikTok 977,896,256 93,358 0.00001
-Snapchat 4,068,586 2,142 0.00053
-Second, we wish to explore how the main result is affected by the volume of illegal content, which varies across platforms and countries.
-To estimate a plausible range for $p$, let us analyze two sources of empirical data.
-Table 1 shows that the fraction of illegal SoRs based on the DSA-TDB ranged from 0.00001–0.00380.
-We can also obtain insights from the Future of Free Speech policy report (28).
-The authors analyzed comment data collected from Facebook and YouTube across Germany, France, and Sweden during two weeks in June–July, 2023.
-They report the fraction of comments that disappear ($\frac{\text{disappear}}{\text{total}}$) and the proportion of disappeared comments that are legal ($\frac{\text{legal}}{\text{disappear}}$).
-From this, we estimate the fraction of comments that are illegal: $\frac{\text{illegal}}{\text{total}}=\frac{\text{disappear}}{\text{total}}\times\frac{\text{illegal}}{\text{disappear}}=\frac{\text{disappear}}{\text{total}}\times(1-\frac{\text{legal}}{\text{disappear}})$.
-Table 2 shows that this apprach yields a range between 0.00002–0.009.
 Table 1: Data from the Future of Free Speech policy report and derived ratios of illegal content.
@@
 Sweden (YT) 0.041 0.946 0.002
-Based on these estimates, we run experiments with $10^{-4}\leq p\leq 10^{-2}$ (Values $p<10^{-4}$ would require a network with infeasibly large size, as discussed in Methods.) In our main analysis we report the reduction in illegal content prevalence, $1-\frac{\bar{I}_{\text{removal}}}{\langle\bar{I}_{\text{baseline}}\rangle}$.
+Second, we wish to explore how the main result is affected by the volume of illegal content, which varies across platforms and countries.
+To estimate a plausible range for $p$, let us analyze the Future of Free Speech policy report (34).
+The report includes comment data collected from Facebook and YouTube across Germany, France, and Sweden during two weeks in June–July 2023.
+It also includes the fraction of comments that disappear ($\frac{\text{disappear}}{\text{total}}$) and the proportion of disappeared comments that are legal ($\frac{\text{legal}}{\text{disappear}}$).
+From this, we estimate the fraction of comments that are illegal: $\frac{\text{illegal}}{\text{total}}=\frac{\text{disappear}}{\text{total}}\times\frac{\text{illegal}}{\text{disappear}}=\frac{\text{disappear}}{\text{total}}\times(1-\frac{\text{legal}}{\text{disappear}})$.
+Table 1 shows that this apprach yields a range between 0.00002–0.009.
+We run experiments using these estimates (except for the smallest value, which would require a network with infeasibly large size, as discussed in Methods).
+In our main analysis we report the reduction in illegal content prevalence, $1-\frac{\bar{I}_{\text{removal}}}{\langle\bar{I}_{\text{baseline}}\rangle}$.
 To show that this ratio is independent of $p$, we plot the scaling relationship between $\bar{I}_{\text{removal}}$ and $\bar{I}_{\text{baseline}}$ in Fig.
@@
 5b the reduction in the prevalence of illegal content for different values of $p$ from Table 1.
-None of the differences between the values in our main results ($p=0.01$) and those for smaller $p$ values are statistically significant.
-These results suggest that our main results are robust with respect to the volume of illegal content in the system.
+We observe more noise for small $p$, but none of the differences between the values in our main results ($p=0.01$) and those for smaller $p$ values are statistically significant.
+These analyses suggest that our main results are robust with respect to the volume of illegal content in the system.
 Figure 5: Robustness analysis for illegal content probability.
@@
 Each data point represents the proportion of illegal content in the system with and without removal.
-Results are based on 40 simulations for each value of $p$ from Table 2, using $\tau=2$.
+Results are based on 40 simulations using $\tau=2$.
+Each simulation is done with an empirical estimate of $p$ (vertical lines) for Facebook and YouTube users from Germany, France, and Sweden, derived from the data reported in Table 1.
+Estimates of $p$ are based on the proportion of disappeared content that is subsequently judged to be legal.
 The gray line represents a least-squares fit of the log-transformed data, yielding a slope of $0.94\pm 0.01$.
-The vertical lines mark the illegal content probabilities from Table 2.
 (b) Illegal content reduction for different illegal content probabilities corresponding to the same countries and platforms, based on 10 simulations with $\tau=2$ (top) and $\tau=8$ (bottom).
@@
 Discussion
-This paper examines empirical data on illegal content takedown on multiple social media platforms.
-Our agent-based model calibrated on this data reveals that removal of illegal content within a one-day window significantly reduces its prevalence and exposure, whereas more prolonged delays allow illegal content to permeate user feeds, diminishing the impact of the moderation.
-Delays beyond 46 days have no impact on the spread of illegal content.
+This paper reveals how delays in illegal content takedown impact its effectiveness.
+Using empirical takedown data for five social media platforms from the DSA-TDB, we find large variations in average takedown delays.
+While platforms like TikTok typically respond within a week, illegal content often remains online for over a month on platforms like Facebook, Instagram, Snapchat, and YouTube.
+Our agent-based model calibrated on this data shows that long delays can substantially diminish the impact of content removal.
+Although this relationship may appear intuitive, our analysis reveals non-linear dynamics between takedown speed and moderation effectiveness, including critical delay thresholds.
+Specifically, we show that removing illegal content within one day significantly reduces its prevalence and exposure, whereas prolonged delays allow illegal content to permeate user feeds — beyond 46 days, takedown becomes futile in curbing the spread of illegal content.
 These findings have practical implications for social media moderation and regulation.
 By adopting more proactive takedown policies, platforms may reduce moderation delay and limit the spread of illegal content.
-Strengthening the use of and cooperation with trusted flaggers, who identify and report content violating community guidelines or legal standards, can further enhance content moderation.
+Another approach to improving content moderation is to strengthen collaboration with trusted flaggers — individuals or organizations that identify and report content violating community guidelines or legal standards.
 With human flaggers being a limited resource (35), efforts should be directed towards prioritizing the most time-sensitive cases.
-For legislation, the importance of prompt and timely moderation appears to support the argument for implementing tight illegal content removal deadlines, as imposed under NetzDG in Germany and KoPl-G in Austria prior to the DSA.
+For legislation, the importance of timely moderation appears to support the argument for implementing tight illegal content removal deadlines, as imposed under NetzDG in Germany and KoPl-G in Austria prior to the DSA.
 However, the practical impact of such deadlines should be carefully assessed in regulatory debates for two reasons.
 First, regulatory takedown deadlines can only address the timespan between the notification of illegal content and its removal, not the period during which illegal content circulates in the network undetected.
-As a result, the impact of tight deadlines on the spread of illegal content is inherently limited.
+Therefore, the impact of tight deadlines on the spread of illegal content is inherently limited.
 In fact, most platforms already report fast reactions after notifications: a median time below 24 hours according to 2024 DSA Transparency Reports and removal of 80% of illegal content within 24 hours according to 2022 national Transparency Reports for Germany and Austria.
-Second, tight takedown deadlines may reduce the time necessary for properly assessing the (un)lawfulness of content, increasing the risk of false moderation decisions and lowering the quality of content moderation.
+These figures suggest that a large portion of the long delays measured from the DSA-TDB data occurs before illegal content is detected or flagged.
+To better inform policy, platforms should be required to report not only the timestamps for publication and removal of content, but also the detection/notification timestamps to the DSA Transparency Database.
+This would allow to identify how long content has circulated in the network undetected as well as the moderation delay in response to notifications.
+Second, tight takedown deadlines may reduce the time necessary for properly assessing the (un)lawfulness of content, increasing the risk of wrong moderation decisions and thus lowering the quality of content moderation.
 Flagged content requires careful verification, including report reviews and appeal resolutions, before removal decisions can be made.
 Any regulatory framework must ensure effective moderation of illegal content while remaining practical for platforms, and mindful of potential adverse effects.
-For example, in France, the Loi Avia was struck down by the Constitutional Council (30) due to concerns about restrictions on freedom of expression (31, 32).
+For example, in France, the Loi Avia was struck down by the Constitutional Council (36) due to concerns about restrictions on freedom of expression (22, 23).
 These considerations suggest that platforms should remove immediately only obviously illegal content.
@@
 Until the assessment is completed, platforms may implement alternative interventions to reduce the circulation of potentially illegal content.
-Our model has several limitations.
-Certain aspects, such as the user activity levels, are calibrated on data from global Twitter users, while the DSA data focuses on European users.
-Unknown biases could stem from this combination of data sources.
-The model further presumes uniform spreading dynamics across various types of illegal content.
-In reality, the appeal of such content may vary widely; scams may attract a broader audience while human trafficking posts may be aimed at specific communities.
-Due to the lack of data on such content, our model does not distinguish among classes of illegal content or between illegal and legal content.
-If illegal content were more appealing, our findings would underestimate the adverse effects of time delay on moderation.
-For modeling content moderation, we make the simplifying assumption that when illegal content is removed, it is deleted from all user feeds.
-In practice, however, some content can be illegal only in certain jurisdictions and access can be selectively disabled for users in these countries.
+While our findings are robust, several limitations warrant consideration.
+The lack of empirical data on illegal content results in several assumptions about the content production process in our model.
+First, the model focuses on the spread of content already posted, without accounting for the user incentives or legal risks influencing decisions to post illegal material.
+We assume static probability distributions of sharing such content, though in reality, this behavior may evolve in response to moderation policies, potentially declining under stricter or faster moderation.
+Second, as mentioned in Methods, we assume uniform appeal and spreading potential across all types of content.
+If illegal content is generally more engaging, the model may underestimate the impact of delayed takedown, and vice versa.
+Lastly, we triangulate data from multiple sources, which may introduce inconsistencies.
+For instance, user activity is calibrated using global Twitter data, whereas the DSA database focuses on European users.
+This mismatch could lead to unknown biases in our calibration.
+Our model of the moderation process also has limitations.
+The DSA-TDB data only captures the downstream outcomes of content moderation, neglecting the detection process.
+Our model thus treats all illegal content as equally detectable, which may overlook important variation in the detection of different content types.
+We also make the simplifying assumption that when illegal content is removed, it is deleted from all user feeds.
+In practice, however, some content may be illegal only in certain jurisdictions and access may be selectively disabled for users in these countries.
 Our model does not consider such a differentiation.
-The DSA-TDB has its own limitations.
-The database depends heavily on platform voluntary reporting, leading to inconsistencies across platforms.
-This reflects a fundamental weakness in self-reporting systems, as moderation decisions are often shaped by specific platform terms of service and community guidelines.
-Future research could extend our model by accounting for the fact that users often alter their posting patterns in response to moderation decisions (33, 34).
-The model could also compare the effects of takedown versus alternative types of interventions, as well as explore combinations of different intervention instruments.
+The database itself has limitations.
+It relies heavily on voluntary reporting by platforms, leading to inconsistent coverage and reporting quality both across and within platforms (31).
+This highlights a broader issue with self-reporting systems.
+Moderation decisions are shaped by platform-specific policies and guidelines, which may not always align with legal standards: content labeled as “illegal” may in fact be lawful but removed due to internal platform rules.
+The issue of false positives may have inflated the inferred probability of illegal content from the DSA-TDB.
+Future research could extend our model by accounting for the fact that users often alter their posting patterns in response to moderation decisions (37, 38).
+One could also extend the model to compare the effects of takedown versus alternative types of interventions, as well as explore combinations of different intervention instruments.
 Finally, future models should consider the trade-offs between the harm caused by persistent illegal content and the unitended effects of moderation tools, e.g., the limitation of free speech caused by false-positive errors.
-Methods
-Model extensions
-We extend SimSoM to model the heterogeneous activity of social media users, where most agents have low activity levels, and only a few are hyperactive.
-We model this by assigning an activity level $a$ to each user at the beginning of the simulation.
-We draw this activity from a power law $P(a)\sim a^{\gamma}$, where we obtain the exponent $\gamma=2.85$ by fitting empirical measurements of average daily posts (Fig.
-1a).
-If $a<1$, the user (re)shares a single message with probability $a$ at each time step; else the user (re)shares $\lfloor a\rfloor$ messages.
-We wish to model the probability of posting illegal messages by low- and high-risk accounts.
-Given the lack of empirical data about this probability, we used the posting of content from low-credibility sources as a proxy.
-Leveraging news source credibility ratings from NewsGuard (35), we categorized a sample of over 250 thousand Twitter users into a low-credibility and a high-credibility group by comparing the average ratings of sources they shared against two thresholds.
-We then inspected the distributions for posting low-credibility content by these two groups and found that both could be modeled using the parametric family of beta distributions.
-Based on this proxy, we model the probability $i$ of posting illegal messages by users in each group (low- and high-risk) with a beta distribution $P(i)\sim i^{\alpha-1}(1-i)^{\beta-1}$, where $\alpha>0$ and $\beta>0$ are parameters that express the mean $\frac{\alpha}{\alpha+\beta}$ and skewness of the distribution for that group.
-Given the two group distributions, we can express the average illegal content probability across the full population as $p=s_{H}\frac{\alpha_{H}}{\alpha_{H}+\beta_{H}}+(1-s_{H})\frac{\alpha_{L}}{\alpha_{L}+\beta_{L}}$, where $s_{H}$ is the relative size of the high-risk groups, and the $\alpha$ and $\beta$ subscripts indicate the groups.
-Since platforms typically do not provide access to original content, there is insufficient scrutiny to determine the legality of content and limited data on the share of illegal content (for an exception see Wagner et al.
-(36)).
-Hence, it is difficult to empirically calibrate the illegal content probability $p$ by estimating the percentage of illegal messages in the system.
-Therefore, we choose different values of $\alpha$ and $\beta$ to explore a broad range of values $10^{-4}\leq p\leq 10^{-2}$.
-Content removal
-The illegal content prevalence is defined as the average ratio of illegal content in circulation at time $t$: $I_{t}=\frac{1}{N}\sum_{i=1}^{N}f_{i,t}$ where $f_{i,t}$ is the fraction of illegal content in user $i$’s feed at time $t$ and $N$ is the number of agents.
-The simulation ends once the system reaches a steady state, signaled by a convergence in the illegal content prevalence.
-To this end, we calculate an exponential moving average at time $t$ as $\bar{I}_{t}=\rho\bar{I}_{t-1}+(1-\rho)I_{t}$, where $\rho$ is a parameter regulating the importance of older values.
-We stop the simulation when the difference between two consecutive values of this moving average is smaller than a threshold, $|\bar{I}_{t}-\bar{I}_{t-1}|/\bar{I}_{t-1}<\epsilon$.
-The parameters $\rho=0.9$ and $\epsilon=0.0001$ used in the reported simulations were tested to ensure that the system’s illegal content prevalence stabilizes at the steady state.
-For each set of parameters, we run 70 simulations without any content removal; these are averaged to obtain baseline metrics.
-We then run 70 simulations with content removal; the reported reduction is averaged across these runs.
-For example, the reduction in illegal content prevalence is calculated by $(\langle\bar{I}_{\text{baseline}}\rangle-\bar{I}_{\text{removal}})/\langle\bar{I}_{\text{baseline}}\rangle$, where $\bar{I}$ is the illegal content prevalence at the end of a simulation and $\langle\bar{I}_{\text{baseline}}\rangle$ is the baseline average.
-All simulations start from random conditions.
-Data sources
-Under the DSA (Art.
-15), platforms must publish Transparency Reports detailing “the median time needed for taking the action.” While platforms comply, reporting practices differ significantly, preventing thorough comparison among platforms.
-Moreover, median turnaround times in transparency reports only cover the time between notification and removal of content, not the time when problematic content circulates in the network before notification.
-Finally, DSA transparency reports do not indicate moderation delays specific to illegal content, and platforms disallow external access to verify data quality.
-More detailed data, suited for our purpose, is available in the newly introduced DSA Transparency Database, which compiles Statements of Reasons (SoRs) explaining single moderation decisions to affected users.
-These SoRs include the reasons for moderation, as well as information on dates of posting and moderation decisions.
-We therefore use this data source to estimate the empirical half-life ($\tau$) values.
-We collect SoRs from the DSA-TDB between January 01 and December 31, 2024.
-We capture relevant records using the strictest filter, where the platforms explicitly state that the moderation is due to illegal content.
-[^1] We select SoRs that have DECISION_GROUND_ILLEGAL_CONTENT in the decision_ground field.
-The database also includes the value DECISION_GROUND_INCOMPATIBLE_CONTENT for the same field.
-This can cover content that is illegal, which would be indicated in an optional boolean field incompatible_content_illegal.
-Since platforms do not use this reporting category consistently, we did not include it in our dataset for calibration.
-We consider four major social media platforms with a sufficient number of relevant SoRs: Facebook, Instagram, YouTube, TikTok, and Snapchat.
-We exclude LinkedIn and Pinterest due to a lack of SoRs related to illegal content and Twitter/X due to suspicious patterns revealed in its DSA data, as also highlighted in prior work (37).
-This results in 8,613 SoRs submitted by Facebook, 3,005 by Instagram, 408,400 by YouTube, 93,358 for TikTok, and 2,142 by Snapchat.
-The time delay in illegal content removal is defined as the difference between the moderation date (application_date field) and the content posting date (content_date field).
-Follower network
-Simulations are conducted on a follower network derived from empirical Twitter data.
-The network connects users obtained from a 10% random sample of public tweets (38).
-The dataset comprises users who posted links to news articles, with likely automated accounts excluded.
-The original follower network constructed from this dataset has $58{,}296$ nodes and $10{,}632{,}023$ edges.
-We reduce the size of this network to speed up the simulations in our main experiment (Section Effects of time delay).
-We employ $k$-core decomposition to identify $N=10{,}006$ nodes that constitute the $k=94$ core.
-Subsequently, a random selection of edges is removed to match the density of the original network.
-This procedure produces $E=1{,}809{,}798$ edges, with each node having an average of roughly 180 friends or followers.
-For the robustness analyses varying high-risk group size and illegal content probability (Section Robustness), we need a large follower network to run meaningful scenarios using low values of these parameters ($s_{H}=10^{-3}$ and $p=10^{-4}$).
-Therefore we use the original network, but remove a random set of edges to speed up the simulations.
-The average in/out-degree is thus reduced to 18 friends per node, resulting in a network with $N=58{,}296$ nodes and $E=1{,}063{,}202$ edges.
-To test the robustness of our results with respect to different network structures, we create a third, synthetic network with features ubiquitous in real-world networks, namely, the presence of hubs and clustering (directed triads).
-To this end, we use a directed variant of the random-walk growth model (39).
-We configure the parameters of this model to generate a network comparable in size and average degree to the network used in the main experiment.
-In particular, we initialize the network with 181 fully connected nodes and assume new nodes have fixed out-degree $k_{out}=180$.
-New nodes are added by first randomly selecting one node to follow.
-For the remaining $k_{out}-1$ connections, each new node has an equal probability to follow a random friend of the initially chosen node or a randomly selected node in the network.
-This results in a network with $N=10{,}006$ nodes and $E=1{,}810{,}905$ edges; each node has on average approximately 180 friends/followers.
 Author contributions

### v3

@@
-Delayed takedown of illegal content on social media makes moderation ineffective
+Audit of takedown delays across social media reveals failure to reduce exposure to illegal content
 Abstract
@@
 However, the impact of the speed of content removal on prevalence, reach, and exposure to illegal content remains underexplored.
-This study examines the relationship with a systematic analysis of takedown delays using data from the EU Digital Services Act Transparency Database, covering five major platforms over a one-year period.
+This study examines the relationship with a systematic audit of takedown delays using data from the EU Digital Services Act Transparency Database, covering five major platforms over a one-year period.
 We find substantial variation in takedown delay, with some content remaining online for weeks or even months.
 To evaluate how these delays affect the prevalence and reach of illegal content and exposure to it, we develop an agent-based model and calibrate it to empirical data.
-We simulate illegal content diffusion, revealing that rapid takedown (within hours) significantly reduces prevalence, reach, and exposure to illegal content, while longer delays fail to reduce its spread.
-Though the effect of delay may seem intuitive, our simulations quantify exactly how takedown speed shapes the spread of illegal content.
+We simulate illegal content diffusion, revealing that rapid takedown (within hours) significantly reduces prevalence, reach, and exposure to illegal content, while the longer delays measured by the audit fail to reduce its spread.
+Though the link between delay and spread is intuitive, our simulations quantify exactly how takedown speed shapes exposure to illegal content.
 Building on these results, we point to the benefits of faster content removal to effectively curb the spread of illegal content, while also considering the limitations of strict enforcement policies.
 Introduction
-Illegal and harmful content — including copyright violations, child sexual abuse material (CSAM), and incitement to violence or terrorism — poses serious harm and demands effective moderation efforts (1, 2).
+Illegal and harmful content — including copyright violations, child sexual abuse material (CSAM), and incitement to violence or terrorism — poses serious harm and demands effective moderation efforts Yar (2018); Jain et al.
+(2020).
 Platforms have developed a broad range of moderation tools and strategies that vary in severity.
-For some problematic content, such as misinformation and hate speech, interventions have been designed to promote critical user engagement, such as sensitive content labels (3, 4), accuracy reminders (5, 6, 7, 8), and friction mechanisms in content sharing (9, 10, 11).
+For some problematic content, such as misinformation and hate speech, interventions have been designed to promote critical user engagement, such as sensitive content labels Morrow et al.
+(2022); Martel and Rand (2023), accuracy reminders Bhuiyan et al.
+(2018); Pennycook et al.
+(2020, 2021); Pennycook and Rand (2022), and friction mechanisms in content sharing Jahanbakhsh et al.
+(2021); Jahn et al.
+(2023); Tomalin (2023).
 These “soft” interventions alter the environment to nudge user behaviors and are thus contingent on user motivations.
-Other techniques reduce the visibility of problematic content by limiting its searchability, recommendation, and engagement (12, 13, 14).
-Stricter interventions include content removal (takedown), account suspensions, and permanent bans (15, 16, 17).
-These approaches enable swift disruption of content virality and minimize exposure, but also raise concerns about censorship (18).
-Given the trade-offs of different moderation approaches, previous modeling work has shown that combining interventions is an effective approach to leverage the strengths of each while mitigating their limitations (19).
+Other techniques reduce the visibility of problematic content by limiting its searchability, recommendation, and engagement Gillespie (2022); Macdonald and Vaughan (2024); Saurwein et al.
+(2025).
+Stricter interventions include content removal (takedown), account suspensions, and permanent bans Chandrasekharan et al.
+(2017); Ali et al.
+(2021); Jhaver et al.
+(2021).
+These approaches enable swift disruption of content virality and minimize exposure, but also raise concerns about censorship Jiang et al.
+(2023).
+Given the trade-offs of different moderation approaches, previous modeling work has shown that combining interventions is an effective approach to leverage the strengths of each while mitigating their limitations Bak-Coleman et al.
+(2022).
 Nevertheless, for clear-cut cases of illegal content, removal or deactivation in concerned jurisdictions remains a legal demand for social media platforms.
@@
 Laws such as Germany’s Network Enforcement Act (NetzDG), Austria’s Communication Platform Act (KoPl-G), and France’s Loi Avia required platforms to remove “obviously illegal” content within 24 hours and other illegal content within seven days.
-These provisions reflected a clear regulatory emphasis on speed and have drawn criticism both for not being strict enough and for being too strict — potentially infringing on free expression (20, 21, 22, 23).
+These provisions reflected a clear regulatory emphasis on speed and have drawn criticism both for not being strict enough and for being too strict — potentially infringing on free expression Kozyreva et al.
+(2023); Schulz et al.
+(2019); Heldt (2020); Vie publique (2020).
 One of the critical issues is the trade-off between moderation speed and accuracy: removing content very quickly can reduce its spread but may also lower precision, causing false-positive errors and legitimate concerns about over-moderation.
@@
 As legal requirements have shifted, the question of how takedown speed affects the prevalence, reach, and exposure of illegal content becomes highly relevant.
-However, thus far, policies enforcing takedown speed lack empirical grounding, making it difficult for policymakers to design and evaluate appropriate policies.
+However, thus far, there is no empirical data on the effects of takedown speed, making it difficult for policymakers to design and evaluate appropriate policies.
 This paper provides a comprehensive evaluation of the effects of moderation speed by proposing a model that measures how prevalence, exposure, and reach of illegal content change in response to takedown delay.
-In addition, we scrutinize available data on takedown delay of selected platforms and discuss the consequences and limitations of regulatory takedown deadlines.
-Available research on the speed of content removal shows that delay in response reduces prevalence and exposure (24, 25).
+In addition, we audit available data on takedown delay of selected platforms and discuss the consequences and limitations of regulatory takedown deadlines.
+Available research on the speed of content removal shows that delay in response reduces the moderation effectiveness on prevalence and exposure Schneider and Rizoiu (2023); Goldstein et al.
+(2023); Edelson et al.
+(2025).
 However, the existing studies largely focus on mis- and disinformation, such as misleading claims or conspiracy theories, which may not generalize to illegal content.
-Moreover, much of this work centers on emergencies, such as COVID-19 anti-vaccination narratives (26) or the January 6 Capitol Riot (25), during which platforms adopted exceptional measures.
+Moreover, much of this work centers on emergencies, such as COVID-19 anti-vaccination narratives Broniatowski et al.
+(2022) or the January 6 Capitol Riot Goldstein et al.
+(2023), during which platforms adopted exceptional measures.
 To inform legal frameworks, it is necessary to examine the standard operating procedures of platforms during routine, non-emergency periods.
-Studying the impact of content removal policies at scale is challenging, both because it would be ethically impermissible to experiment with harmful or illegal content, and because data on removals is often inaccessible due to privacy concerns.
+Studying the impact of content removal policies at scale is challenging, both because it would be ethically impermissible to experiment with harmful or illegal content, and because provision of access to illegal content is critical and involves privacy concerns.
 Currently, no studies systematically examine illegal content removal practices across platforms or quantify the effects of takedown delays.
-Based on a year of data from the DSA Transparency Database (DSA-TDB) (27), we show that illegal content removal speed varies widely across platforms, ranging from several days to weeks, and in some cases, up to a year.
+Based on a year of data from the DSA Transparency Database (DSA-TDB) Commission (2024), we show that illegal content removal speed varies widely across platforms, ranging from several days to weeks, and in some cases, up to a year.
 By calibrating an agent-based model with empirically estimated takedown delays, we quantify the relationship between moderation speed and the reduction of illegal content spread.
@@
 However, in the Discussion, we point to limitations of takedown deadlines and argue that any legal framework should balance the goal of timely removal with the need to ensure the quality of moderation and account for platform-level operational constraints.
+Figure 1: Pipeline for estimating the effectiveness of illegal content removal.
+(a) Empirical data calibration: We estimate a plausible range for the probability of illegal content $p$ from policy reports and infer platform-specific takedown delays from Statements of Reasons (SoRs) in the DSA Transparency Database (DSA-TDB) for five major platforms: Facebook, Instagram, YouTube, TikTok and Snapchat.
+These delays are summarized by the expected content takedown delay $\tau$ and used to parameterize content removal.
+Simulations are run on an empirical follower network derived from Twitter data.
+(b) Information diffusion and removal: An agent-based model simulates the spread of content over the network.
+At each timestep, illegal content survives with constant probability $p_{s}$, calibrated from the empirical takedown delay $\tau$; once removed, content is deleted from all user feeds.
+(c) Moderation effectiveness estimation: Simulations are performed across a broad range of takedown delays to quantify the impact of removal speed.
+The effect of content removal is measured as the relative reduction in the prevalence and exposure of illegal content, compared with a baseline without removal.
 Methods
-Modeling social media diffusion
-We use SimSoM (28), an agent-based model that mimics information diffusion on social networks such as Twitter/X, Mastodon, Bluesky, Threads, or Instagram.
-The platform is represented as a directed network, where nodes correspond to user accounts and edges denote follower relationships, which need not be reciprocal.
-As in real-world platforms, information circulates through messages that appear on user news feeds.
-At each time step, a user can post new original messages, reshare content from their feeds, or take no action.
-Messages represent information that could take the form of text, links, hashtags, images, or other media.
-To select a message for resharing, an agent samples from the news feed, which contains distinct messages shared by their friends.
-Once shared by an agent, a message appears in the news feeds of their followers, allowing for propagation across the network.
-Thus, the exposure of the message depends on its author’s follower count — agents with more followers will achieve greater exposure.
-Capturing mechanisms of real-world ranking algorithms, agents prefer to reshare messages that are appealing, recent, and popular (29).
-The appeal of a message is a scalar value that captures the intrinsic attractiveness of messages.
-This could derive from properties such as novelty or emotional appeal.
-The distribution of message appeal captures a broad distribution of virality often seen in real-world information diffusion networks (28).
-The recency of a message decreases as time passes since it was first introduced to the agent’s news feed.
-The popularity of a message is the number times it was reshared by the agent’s friends.
-Building on the base model, we incorporate additional features to examine the diffusion of illegal content.
-Specifically, each message includes a binary legality attribute indicating whether the content is legal or illegal.
-Different types of illegal content may be more or less appealing.
-For example, CSAM is criminally and morally condemned, therefore resharing is expected to be rare.
-More common forms of illegal content, such as copyrighted media, tend to be highly engaging.
-Lacking data on the correlation between legality and appeal, we make the simplifying assumption that the two are independent.
-Figure 1: User activity distribution and probability of posting illegal content.
-(a) Complementary cumulative distribution (CCDF) of daily average Twitter user activity based on 81,633,118 posts by 37,882,999 accounts during nine days in January 2023 (orange line).
-The dashed line provides a scaling reference, with the slope 2.85-2.85 representing the exponent of the power law fit.
-(b) Illegal content posting distributions for low-risk (orange) and high-risk (purple) users, for mean illegal content probability $p=10^{-4}$ (top), $p=10^{-3}$ (middle), and $p=10^{-2}$ (bottom), based on a large random sample.
-We also extend SimSoM to model the heterogeneous activity of social media users, where most agents have low activity levels, and only a few are hyperactive.
-We model this by assigning an activity level $a$ to each user at the beginning of the simulation.
-We draw this activity from a power law $P(a)\sim a^{\gamma}$, where we obtain the exponent $\gamma=2.85$ by fitting empirical measurements of average daily posts (Fig.
-1a).
-If $a<1$, the user (re)shares a single message with probability $a$ at each time step; else the user (re)shares $\lfloor a\rfloor$ messages.
-In the following sections, we outline our approach to examine the relationship between takedown delay and effectiveness.
-We first describe the modeling of illegal content diffusion and takedown, followed by an analysis of the DSA-TDB to estimate expected takedown times, and finally, we detail the experimental setup.
-Modeling illegal content diffusion and removal
-We assume that illegal content can be generated by any user, though only a small fraction $s_{H}$ of “high-risk accounts” (H) frequently share such content — for example, CSAM, scams, or terrorist propaganda.
-The rest of the network, “low-risk accounts” (L) may occasionally violate the law, such as sharing copyrighted content.
-The illegal content probability $p$ represents the aggregated probability across user groups, capturing the average likelihood that a user in the population posts illegal content.
-Specifically, the probability of posting illegal messages by both low- and high-risk accounts follows a beta distribution $P(i)\sim i^{\alpha-1}(1-i)^{\beta-1}$, where $\alpha>0$ and $\beta>0$ are parameters that express the mean $\frac{\alpha}{\alpha+\beta}$ and the skewness of the distribution for that group.
-Given the two group distributions, we can express the average illegal content probability across the full population as $p=s_{H}\frac{\alpha_{H}}{\alpha_{H}+\beta_{H}}+(1-s_{H})\frac{\alpha_{L}}{\alpha_{L}+\beta_{L}}$, where $s_{H}$ is the relative size of the high-risk groups, and the $\alpha$ and $\beta$ subscripts indicate the groups.
-Fig.
-1b plots the probability distributions of posting illegal messages by each group for different values of $p$.
-Since platforms typically do not provide access to original content, there is insufficient scrutiny to determine the legality of content and limited data on the share of illegal content (for an exception see Wagner et al.
-(30)).
-Hence, it is difficult to empirically calibrate the illegal content probability $p$ by estimating the percentage of illegal messages in the system.
-In our experiments, unless stated otherwise, we set $s_{H}=0.1$ to reflect a minority of high-risk accounts.
-We use the values of $\alpha_{L}=0.1$, $\beta_{L}=90$ for the low-risk group and $\alpha_{H}=3$, $\beta_{H}=30$ for the high-risk group, resulting in $p=10^{-2}$.
-We show that the results are robust to these parameter choices in the Robustness section.
-To understand the dynamics of illegal content removal, let us consider the lifespan of such content on a platform.
-While some illegal content is immediately detected and removed using automated detection and decision systems, other is missed due to detection limitations.
-In such cases, content moderation is delayed until platforms either detect the content themselves later or receive external notification.
-Additional delays may occur when platforms assess the content against their community standards and national laws.
-If a violation is confirmed, either the content is removed globally or access is restricted in certain jurisdictions.
-We model takedown delay by making the simplifying assumption that at any time step, there is a constant probability $p_{s}$ that a piece of illegal content is not detected and removed, and therefore survives.
-This yields an exponential survival curve for the delay $t$, $P(t)\sim 2^{-t/\tau}$, where the illegal content half-life $\tau$ is the time taken for half of the illegal content to be removed (median removal time since publication).
-We use $\tau$ to parameterize content takedown delay.
-For easier interpretation, we can alternatively define the expected time delay as $\tau/\ln 2$.
-Since $P(t)\approx p_{s}^{t}$ in the simulation, we set $p_{s}=2^{-1/\tau}$.
-When a message is removed, it is deleted from all user feeds.
-In the next section we derive a realistic range of $\tau$ values from empirical data for five major social media platforms.
-DSA data analysis
-Figure 2: Distributions (CCDF) of time delay in illegal content takedown on five platforms.
-Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line).
-Top panels: Fits based on raw data with linear y-axes.
-Bottom panels: Fits based on a linear-log transformation of the data, highlighting extreme time delay values.
-Under the DSA (Art.
-15), platforms must publish Transparency Reports detailing “the median time needed for taking the action.” While platforms comply, reporting practices differ significantly, preventing thorough comparison among platforms.
-Moreover, median turnaround times in transparency reports only cover the time between notification and removal of content, not the time when problematic content circulates in the network before notification.
-Finally, DSA transparency reports do not indicate moderation delays specific to illegal content, and platforms disallow external access to verify data quality.
-More detailed data, suited for our purpose, is available in the DSA Transparency Database, which compiles Statements of Reasons (SoRs) explaining single moderation decisions to affected users.
-These SoRs include the reasons for moderation, as well as information on dates of posting and moderation decisions.
-We therefore use this data source to estimate the empirical half-life ($\tau$) values.
+We begin by analyzing the DSA-TDB to understand the illegal content removal practice on different platforms.
+We then introduce a three-step pipeline to evaluate the effectiveness of moderation under different time delays using an agent-based model, as outlined in the following sections and illustrated in Fig.
+1.
+We first describe the empirical estimation of the parameters used to calibrate the simulations.
+These include (i) the probability that newly created content is illegal, estimated from external policy reports; and (ii) the expected takedown delays of illegal content, derived from the DSA-TDB.
+Next, we describe the extensions made to SimSoM, an agent-based model of social media, to capture information diffusion and content removal mechanisms.
+Finally, we detail the experimental setup used to run multiple simulation scenarios.
+Estimating takedown delay from the DSA Transparency Database
+To study cross-platform takedown speed for illegal content, we rely on the DSA Transparency Database (DSA-TDB) Commission (2024), which provides case-level Statements of Reasons (SoRs) with timestamps for content publication and removal, allowing us to measure how long content remained online.
+Other available takedown data come from platform Transparency Reports, for example under Art.
+15 of the DSA.
+They typically report aggregated notice-to-action times, which capture responsiveness after notification but not the publication-to-removal duration required to measure how long content was online and potentially reachable by users, which is central to our analysis (see Appendix, Alternative platform compliance data).
 We collect SoRs from the DSA-TDB between January 1 and December 31, 2024.
 We capture relevant records using the strictest filter, where the platforms explicitly state that the moderation is due to illegal content.
-[^1] We select SoRs that have DECISION_GROUND_ILLEGAL_CONTENT in the decision_ground field.
-The database also includes the value DECISION_GROUND_INCOMPATIBLE_CONTENT for the same field.
-This can cover content that is illegal, which would be indicated in an optional boolean field incompatible_content_illegal.
+[^1] The DSA-TDB contains reporting fields to indicate the decision ground for content moderation.
+Content must be classified either as illegal (DECISION GROUND ILLEGAL CONTENT), or as incompatible with terms and conditions (DECISION GROUND INCOMPATIBLE CONTENT).
+For our analysis, we select SoRs that classified content as illegal.
+However, since platforms do not have do conduct legal assessments of all moderated content, more than 99% of content is classified as incompatible with terms and conditions, and the share of illegal content may be underreported Trujillo et al.
+(2025); Groesch et al.
+(2026).
+In addition, platforms can indicate in an optional boolean field if content, which was classified as incompatible with terms and conditions, is also considered illegal (INCOMPATIBLE CONTENT ILLEGAL).
 Since platforms do not use this reporting category consistently, we did not include it in our dataset for calibration.
 We consider five major social media platforms with a sufficient number of relevant SoRs: Facebook, Instagram, YouTube, TikTok, and Snapchat.
-We exclude LinkedIn and Pinterest due to a lack of SoRs related to illegal content and Twitter/X due to suspicious patterns revealed in its DSA data, as also highlighted in prior work (31).
-This results in 8,613 SoRs submitted by Facebook, 3,005 by Instagram, 408,400 by YouTube, 90,409 by TikTok, [^2] We excluded 2,949 observations with invalid input for the content creation date (2000-01-01).
-and 2,142 by Snapchat.
+We exclude LinkedIn and Pinterest due to a lack of SoRs related to illegal content and Twitter/X due to suspicious patterns revealed in its DSA data, as also highlighted in prior work Trujillo et al.
+(2025); Groesch et al.
+(2026).
+This results in 8,613 SoRs submitted by Facebook, 3,005 by Instagram, 408,400 by YouTube, 2,142 by Snapchat, and 90,409 by TikTok.
+[^2] We excluded 2,949 TikTok observations with invalid input for the content creation date (2000-01-01).
 The takedown delay is defined as the difference between the moderation date (application_date field) and the content posting date (content_date field).
-Fig.
-2 shows empirical survival time curves for the five social media platforms from DSA-TDB SoR records.
-By fitting exponential survival time curves $P(t)\sim 2^{-t/\tau}$ to this empirical data, we estimate that the half-life is $\tau\approx 21$ days for Facebook, 60 days for Instagram, 198 days for YouTube, 4 days for TikTok, and 94 days for Snapchat.
-The exponential survival curves (Fig.
-2 top) fit the data reasonably well for short delays, less for extended times.
-We can get an intuition for this by noting that the longer the survival of an illegal piece of content, the lower its visibility, and therefore the lower its likelihood to be reported and removed — $p_{s}$ increases instead of remaining constant.
-To better capture the tails of the data, we also estimate $\tau$ by fitting the linear-log transformed data to $\log_{2}P(t)\sim-t/\tau$.
-This procedure yields longer half-life estimates $\tau\approx 390$ days for Facebook, 340 days for Instagram, 786 days for YouTube, 249 days for TikTok, and 360 days for Snapchat (Fig.
-2 bottom).
-While none of the exponential survival curves perfectly fit the empirical takedown delay data, the short and long estimates from these different fits provide reasonable bounds for the half-life parameter.
-We therefore use them to determine a range of average delay for each platform.
-Simulation framework and parameters
-Based on the above analysis, we simulate various takedown scenarios by exploring a broad range of illegal content half-life values, $2^{-4}\leq\tau\leq 2^{9}$ days.
-Illegal content prevalence is defined as the average ratio of illegal content in circulation at time $t$: $I_{t}=\frac{1}{N}\sum_{i=1}^{N}f_{i,t}$ where $f_{i,t}$ is the fraction of illegal content in user $i$’s feed at time $t$ and $N$ is the number of agents.
-The simulation ends once the system reaches a steady state, signaled by a convergence in the illegal content prevalence.
-To this end, we calculate an exponential moving average at time $t$ as $\bar{I}_{t}=\rho\bar{I}_{t-1}+(1-\rho)I_{t}$, where $\rho$ is a parameter regulating the importance of older values.
-We stop the simulation when the difference between two consecutive values of this moving average is smaller than a threshold, $|\bar{I}_{t}-\bar{I}_{t-1}|/\bar{I}_{t-1}<\epsilon$.
-The parameters $\rho=0.9$ and $\epsilon=0.0001$ used in the reported simulations were tested to ensure that the system’s illegal content prevalence stabilizes at the steady state.
-For each set of parameters, we run 70 simulations without any content removal; these are averaged to obtain baseline metrics.
-We then run 70 simulations with content removal; illegal content prevalence and other quantities are averaged across these runs.
-All simulations start from random conditions.
-Simulations are conducted on a follower network derived from empirical Twitter data.
-The network connects users obtained from a 10% random sample of public tweets (32).
-The dataset comprises users who posted links to news articles, with likely automated accounts excluded.
-The original follower network constructed from this dataset has 58,29658{,}296 nodes and 10,632,02310{,}632{,}023 edges.
-We reduce the size of this network to speed up the simulations in our main experiment (Effects of takedown delay section).
-We employ $k$-core decomposition to identify N=10,006N=10{,}006 nodes that constitute the $k=94$ core.
-Subsequently, a random selection of edges is removed to match the density of the original network.
-This procedure produces E=1,809,798E=1{,}809{,}798 edges, with each node having an average of roughly 180 friends or followers.
-For the robustness analyses, we vary high-risk group size and illegal content probability (Robustness section).
-We therefore need a large follower network to run meaningful scenarios using low values of these parameters ($s_{H}=10^{-3}$ and $p=3\times 10^{-4}$).
-Therefore, we use the original network but remove a random set of edges to speed up the simulations.
-The average in/out-degree is thus reduced to 18 friends per node, resulting in a network with N=58,296N=58{,}296 nodes and E=1,063,202E=1{,}063{,}202 edges.
-To test the robustness of our results with respect to different network structures, we create a third, synthetic network with features ubiquitous in real-world networks, namely, the presence of hubs and clustering (directed triads).
-To this end, we use a directed variant of the random-walk growth model (33).
-We configure the parameters of this model to generate a network comparable in size and average degree to the network used in the main experiment.
-In particular, we initialize the network with 181 fully connected nodes and assume new nodes have fixed out-degree $k_{out}=180$.
-New nodes are added by first randomly selecting one node to follow.
-For the remaining $k_{out}-1$ connections, each new node has an equal probability to follow a random friend of the initially chosen node or a randomly selected node in the network.
-This results in a network with N=10,006N=10{,}006 nodes and E=1,810,905E=1{,}810{,}905 edges.
-Results
-Takedown delay
-Our analysis of the DSA-TDB for five major social media platforms reveals substantial variation in illegal content takedown delays — ranging from several days to weeks, and in some cases up to a year or more.
-TikTok has the fastest takedown speed, with most illegal content lasting approximately 6–359 days.
-YouTube has the the slowest content takedown time, where illegal content can take between 286 days and 3 years to be removed.
-The average illegal content life time for other platforms falls between these extremes, also ranging widely: approximately 30–563 days for Facebook, 87–491 days for Instagram, and 136–519 days for Snapchat.
-(See DSA data analysis for how these ranges are obtained.)
-The observed heterogeneity in takedown delays may also reflect differences in the types of illegal content being moderated.
-Some content types, such as copyright violations, can often be verified quickly, while others may require more time for investigation.
-Effects of takedown delay
-We measure the effects of time delay in content takedown by the reduction of illegal content throughout the network, which can be quantified by multiple metrics.
-Prevalence measures the number of posts containing illegal content that are in the system at steady state (see Methods).
-The more illegal content is present in the system, the more harm it can potentially cause.
-However, a piece of illegal content may be removed from a user’s feed before the user sees it.
-Therefore we also measure harm more directly using two exposure metrics.
-Impressions measures the number of times illegal content is seen by users.
-Note that a single user can be exposed to many posts with illegal content, or even the same post multiple times.
-Reach is a complementary measure of exposure, capturing the number of unique users who see any amount of illegal content.
-For each of these measures, we calculate the reduction percentage relative to a baseline measure without any content removal.
-For example, the reduction in illegal content prevalence is calculated by $(\langle\bar{I}_{\text{baseline}}\rangle-\bar{I}_{\text{removal}})/\langle\bar{I}_{\text{baseline}}\rangle$, where $\bar{I}$ is the illegal content prevalence at the end of a simulation and $\langle\bar{I}_{\text{baseline}}\rangle$ is the baseline average.
-Figure 3: Impact of takedown delay on illegal content.
-The reduction in the prevalence (pink), reach (green), and impressions (purple) of illegal content is plotted as a function of the expected takedown delay ($\tau/\ln 2$).
-Shading represents the 95% confidence intervals estimated using non-parametric bootstrapping.
-The vertical lines indicate the expected takedown delays obtained by fitting the DSA data to exponential curves for different platforms.
-Fig.
-3 plots the reduction in illegal content prevalence, impressions, and reach as a function of the expected takedown delay.
-Content removal has maximal impact under rapid takedowns.
-For expected delay below nine hours (corresponding to half-life $\tau<0.25$ days), removal results in approximately 95–100% reduction of illegal content across all measures.
-Moderation becomes less effective as the time delay increases, allowing illegal content to spread before being removed.
-Between approximately 9 hours–46 days ($0.25\leq\tau\leq 32$ days), the reduction in prevalence of illegal content drops from 95% to negligible levels.
-The empirical values measured from DSA-TDB data for TikTok and Facebook are within this interval.
-Those for Instagram, Snapchat, and YouTube correspond to longer delays, where removal has no significant effect on illegal content prevalence.
-The high ends of the delay ranges are even longer, as mentioned in the previous subsection.
-Although the overall trend is similar across the two exposure metrics, the reduction in impressions and reach drops even faster for delays of up to a few days.
-For example, when illegal content is removed within a day on average, reach and impressions are only reduced by about 60%.
-In summary, these findings show that delays beyond a day significantly diminish the effect of content removal.
-Robustness
-The results presented above may depend on the values of $s_{H}$ (proportion of high-risk accounts), $p$ (illegal content probability), and the underlying follower network.
-We therefore investigate if the results remain valid when changing these parameters.
-For these robustness experiments, we use two different values of illegal content half-life, $\tau=2$ and $\tau=8$, and assign users to the high- and low-risk groups randomly in each run.
-First, we run experiments to examine the effect of having smaller high-risk groups ($10^{-3}\leq s_{H}\leq 10^{-1}$), which may better reflect realistic scenarios.
-The main outcome variable, reduction in illegal content prevalence, shows no statistically significant differences (Fig.
-4), suggesting that our findings are robust to changes in $s_{H}$.
-Figure 4: Robustness of results with respect to different high-risk group sizes $s_{H}$ for $\tau=2$ (top) and $\tau=8$ (bottom) based on 20 simulations.
-The illegal content reduction does not vary significantly ($P\geq 0.72$ across pairs of experiments with different $s_{H}$ values, using Mann–Whitney U tests with Bonferroni correction).
+Beyond highlighting variation in content removal practices across platforms, this data source provides a way to empirically estimate the expected takedown delay of illegal content.
+Throughout the paper, we denote this quantity by $\tau$.
+Empirically, $\tau$ summarizes observed moderation behavior in the DSA-TDB; in the simulation, $\tau$ serves as a parameter governing the stochastic removal process.
+Estimating the probability of illegal content
+We define the illegal content probability $p$ as the mean prevalence of illegal content in the system.
+Empirically estimating $p$ is challenging, as platforms rarely provide access to original content or systematic assessments of legality.
+As a result, there is limited direct evidence on the prevalence of illegal content in large-scale online systems (for an exception, see Wagner et al.Wagner et al.
+(2024)).
+To estimate a plausible range for $p$, let us analyze the Future of Free Speech policy report of Free Speech Project (2024).
+The report includes comment data collected from Facebook and YouTube across Germany, France, and Sweden during two weeks in June–July 2023.
+It also includes the fraction of comments that disappear ($\frac{\text{disappear}}{\text{total}}$) and the proportion of disappeared comments that are legal ($\frac{\text{legal}}{\text{disappear}}$).
+From this, we estimate the fraction of comments that are illegal: $\frac{\text{illegal}}{\text{total}}=\frac{\text{disappear}}{\text{total}}\times\frac{\text{illegal}}{\text{disappear}}=\frac{\text{disappear}}{\text{total}}\times(1-\frac{\text{legal}}{\text{disappear}})$.
+This approach treats all disappeared comments as potential removals, regardless of whether deletion was performed by the platform, moderators, or the original author.
+As a result, it may overestimate the true prevalence of illegal content, since some removed content is later classified as legal.
+Nevertheless, it provides a conservative upper bound.
+Table 1 shows that this approach yields a wide range between 0.00002–0.009.
+We run simulations across this range of $p$ values, excluding $p=0.00002$ that would require infeasibly large networks to produce stable estimates.
+We find that system-level metrics are largely independent from the value of $p$ (see Appendix, Robustness); we use $p=0.01$ in the simulations.
+While $p$ captures the mean prevalence of illegal content, it does not determine how such content is distributed across users.
+In practice, illegal content may be generated uniformly across the population or be highly concentrated among a small subset of accounts.
+We examine several alternative distributions of individual-level illegal posting probabilities that all share the same population mean $p$: (i) a near-homogeneous scenario in which all users have similar propensities to post illegal content; (ii) a heterogeneous scenario in which some users have substantially higher probabilities of posting illegal content than others; and (iii) a two-group scenario in which the population consists of a small fraction of high-risk users who repeatedly post illegal content and a majority of low-risk users who only occasionally violate the law (see Appendix, Illegal content probability modeling).
+Results are robust across all three assumptions about how illegal content probability is distributed across users (see Appendix, Robustness).
+We decided to model illegal content as being generated by two groups, each characterized by a beta distribution (see Appendix, Illegal content probability modeling).
 Table 1: Data from the Future of Free Speech policy report and derived ratios of illegal content.
@@
 Sweden (YT) 0.041 0.946 0.002
-Second, we wish to explore how the main result is affected by the volume of illegal content, which varies across platforms and countries.
-To estimate a plausible range for $p$, let us analyze the Future of Free Speech policy report (34).
-The report includes comment data collected from Facebook and YouTube across Germany, France, and Sweden during two weeks in June–July 2023.
-It also includes the fraction of comments that disappear ($\frac{\text{disappear}}{\text{total}}$) and the proportion of disappeared comments that are legal ($\frac{\text{legal}}{\text{disappear}}$).
-From this, we estimate the fraction of comments that are illegal: $\frac{\text{illegal}}{\text{total}}=\frac{\text{disappear}}{\text{total}}\times\frac{\text{illegal}}{\text{disappear}}=\frac{\text{disappear}}{\text{total}}\times(1-\frac{\text{legal}}{\text{disappear}})$.
-Table 1 shows that this apprach yields a range between 0.00002–0.009.
-We run experiments using these estimates (except for the smallest value, which would require a network with infeasibly large size, as discussed in Methods).
-In our main analysis we report the reduction in illegal content prevalence, $1-\frac{\bar{I}_{\text{removal}}}{\langle\bar{I}_{\text{baseline}}\rangle}$.
-To show that this ratio is independent of $p$, we plot the scaling relationship between $\bar{I}_{\text{removal}}$ and $\bar{I}_{\text{baseline}}$ in Fig.
-5a for different $p$ values from Table 1.
-We observe only a weak sublinear relationship, $\bar{I}_{\text{removal}}\sim\bar{I}_{\text{baseline}}^{0.94}$, suggesting that the ratio is nearly independent of $p$.
-We also plot in Fig.
-5b the reduction in the prevalence of illegal content for different values of $p$ from Table 1.
-We observe more noise for small $p$, but none of the differences between the values in our main results ($p=0.01$) and those for smaller $p$ values are statistically significant.
-These analyses suggest that our main results are robust with respect to the volume of illegal content in the system.
-Figure 5: Robustness analysis for illegal content probability.
-(a) Scaling between illegal content prevalence in removal and baseline conditions.
-Each data point represents the proportion of illegal content in the system with and without removal.
-Results are based on 40 simulations using $\tau=2$.
-Each simulation is done with an empirical estimate of $p$ (vertical lines) for Facebook and YouTube users from Germany, France, and Sweden, derived from the data reported in Table 1.
-Estimates of $p$ are based on the proportion of disappeared content that is subsequently judged to be legal.
-The gray line represents a least-squares fit of the log-transformed data, yielding a slope of $0.94\pm 0.01$.
-(b) Illegal content reduction for different illegal content probabilities corresponding to the same countries and platforms, based on 10 simulations with $\tau=2$ (top) and $\tau=8$ (bottom).
-The result does not vary significantly ($P\geq 0.13$ across pairs of experiments with different $p$ values, using Mann–Whitney U tests with Bonferroni correction).
-Last, we examine the effect of network structure by comparing experiments using the empirical network and a synthetic network (see Methods).
-There is no statistically significant difference in the reduction in illegal content prevalence between these two networks, indicating the robustness of our findings (Fig.
-6).
-Figure 6: Robustness analysis with respect to network structures based on 10 simulations with $\tau=2$ (left) and $\tau=8$ (right).
-The illegal content reduction does not vary significantly ($P=0.11$ and $P=0.22$, respectively, using Mann–Whitney U tests with Bonferroni correction).
+Modeling social media diffusion
+We use SimSoM Truong et al.
+(2024), an agent-based model that mimics information diffusion on social networks such as Twitter/X, Mastodon, Bluesky, Threads, or Instagram.
+The platform is represented as a directed network, where nodes correspond to user accounts and edges denote follower relationships, which need not be reciprocal.
+As in real-world platforms, information circulates through messages that appear on user news feeds.
+At each time step, a user can post new original messages, reshare content from their feed, or take no action.
+Messages represent information that could take the form of text, links, hashtags, images, or other media.
+To select a message for resharing, an agent samples from the news feed, which contains distinct messages shared by their friends.
+Once shared by an agent, a message appears in the news feeds of their followers, allowing for propagation across the network (see Discussion).
+Thus, the exposure of the message depends on its author’s follower count — agents with more followers will achieve greater exposure.
+The model captures mechanisms of real-world ranking algorithms and sharing behavior, according to which agents prefer to reshare messages that are appealing, recent, and popular Twitter (2023); Truong et al.
+(2024).
+Building on the base model, we incorporate additional features to examine the diffusion of illegal content.
+Specifically, each message includes a binary attribute indicating whether the content is illegal.
+Different types of illegal content may be more or less appealing.
+For example, CSAM is criminally and morally condemned, therefore resharing is expected to be rare.
+More common forms of illegal content, such as copyrighted media, tend to be highly engaging.
+Lacking data on the correlation between legality and appeal, we make the simplifying assumption that the two are independent.
+We also extend SimSoM to model the heterogeneous activity of social media users, where most agents have low activity levels, and only a few are hyperactive.
+We model this by assigning an activity level $a$ to each user at the beginning of the simulation.
+We draw this activity from a power law $P(a)\sim a^{\gamma}$.
+We estimate the exponent $\gamma=2.85$ by fitting empirical measurements of average daily Twitter activity, based on 81,633,118 posts by 37,882,999 accounts collected over nine days in January 2023.
+If $a<1$, the user (re)shares a single message with probability $a$ at each time step; else the user (re)shares $\lfloor a\rfloor$ messages.
+Modeling illegal content diffusion and removal
+To understand the dynamics of illegal content removal, let us consider the lifespan of such content on a platform.
+While some illegal content is immediately detected and removed using automated detection and decision systems, other is missed due to detection limitations.
+In such cases, content moderation is delayed until the platform either detects the content itself or receives an external notification.
+Additional delays may occur when platforms assess the content against their community standards and national laws.
+If a violation is confirmed, either the content is removed globally or access is restricted in certain jurisdictions.
+While real-world moderation involves multiple detection and decision stages, our objective is to capture the aggregate lifetime during which illegal content can diffuse through the network.
+To this end, we model removal using a stochastic survival process.
+We make the simplifying assumption that at any time step, illegal content survives with probability $p_{s}$ and is removed with probability $1-p_{s}$.
+Under this assumption, the probability that content survives for at least $t$ time steps is $P(t)\approx p_{s}^{t}$, since the content must survive $t$ consecutive removal opportunities.
+Empirically, takedown delays are well described by an exponential survival curve, $P(t)\approx e^{-t/\tau}$, where $\tau$ is the expected takedown delay or mean lifetime obtained by fitting the DSA-TDB data (see Long takedown delays across platforms).
+To ensure that the simulation mirrors the same survival curve and $\tau$ values as observed in the empirical data, we parameterize the discrete-time survival process to match this exponential form at integer time steps.
+Since $P(t)\approx p_{s}^{t}$ in the simulation, we set $p_{s}=e^{-1/\tau}$.
+Thus, at each time step we remove content with probability $1-e^{-1/\tau}$.
+When a message is removed, it is deleted from all user feeds.
+Simulation framework and metrics
+Having specified the diffusion and removal dynamics of illegal content, we now turn to the simulation framework used to evaluate the effects of removal delays — the network structure, outcome metrics, convergence criterion, and parameter settings used in the simulations.
+Simulations are conducted on a follower network derived from empirical Twitter data.
+The network connects users obtained from a 10% random sample of public tweets Nikolov et al.
+(2020).
+The dataset comprises users who posted links to news articles, with likely automated accounts excluded.
+The original follower network constructed from this dataset has 58,29658{,}296 nodes and 10,632,02310{,}632{,}023 edges.
+We reduce the size of this network to speed up the simulations reported in Takedown fails to curb illegal content prevalence and exposure.
+We employ $k$-core decomposition to identify N=10,006N=10{,}006 nodes that constitute the $k=94$ core.
+Subsequently, a random selection of edges is removed to match the density of the original network.
+This procedure produces E=1,809,798E=1{,}809{,}798 edges, with each node having an average of roughly 180 friends or followers.
+The main results are robust with respect to the structure of the network (see Appendix, Robustness).
+We measure the effects of time delay in content takedown by the reduction of illegal content throughout the network, which can be quantified by multiple metrics.
+Illegal content prevalence is defined as the average ratio of illegal content in circulation at time $t$: $I_{t}=\frac{1}{N}\sum_{i=1}^{N}f_{i,t}$ where $f_{i,t}$ is the fraction of illegal content in user $i$’s feed at time $t$ and $N$ is the number of agents.
+The more illegal content is present in the system, the more harm it can potentially cause.
+However, a piece of illegal content may be removed from a user’s feed before the user sees it.
+Therefore we also measure potential impact more directly using two exposure metrics.
+Impressions measures the number of times illegal content is seen by users.
+Note that a single user can be exposed to many posts with illegal content, or even the same post multiple times.
+Reach is a complementary measure of exposure, capturing the number of unique users who see any amount of illegal content.
+For each of these metrics, we calculate the reduction percentage relative to a baseline measure without any content removal.
+For example, the reduction in illegal content prevalence is calculated by $(\langle\bar{I}_{\text{baseline}}\rangle-\bar{I}_{\text{removal}})/\langle\bar{I}_{\text{baseline}}\rangle$, where $\bar{I}$ is the illegal content prevalence at the end of a simulation and $\langle\bar{I}_{\text{baseline}}\rangle$ is the baseline average.
+The simulation ends once the system reaches a steady state, signaled by a convergence in the illegal content prevalence.
+To this end, we calculate an exponential moving average at time $t$ as $\bar{I}_{t}=\rho\bar{I}_{t-1}+(1-\rho)I_{t}$, where $\rho$ is a parameter regulating the importance of older values.
+We stop the simulation when the difference between two consecutive values of this moving average is smaller than a threshold, $|\bar{I}_{t}-\bar{I}_{t-1}|/\bar{I}_{t-1}<\epsilon$.
+The parameters $\rho=0.9$ and $\epsilon=0.0001$ used in the reported simulations were tested to ensure that the system’s illegal content prevalence stabilizes at the steady state.
+For each set of parameters, we run 70 simulations without any content removal; these are averaged to obtain baseline metrics.
+We then run 70 simulations with content removal; illegal content prevalence and exposure metrics (reach and impressions, defined in Takedown fails to curb illegal content prevalence and exposure) are averaged across these runs.
+All simulations start from random conditions.
+Results
+Long takedown delays across platforms
+Figure 2: Complementary cumulative distribution functions (CCDFs) of time delay in illegal content takedown on five platforms.
+Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line).
+Our analysis of the DSA-TDB for five major social media platforms reveals substantial variation in illegal content takedown delays, ranging from several days to weeks and in some cases up to several years.
+Figure 2 presents the empirical distributions of takedown delays derived from DSA-TDB Statements of Reasons for the five platforms.
+By fitting exponential survival models of the form $P(t)\sim e^{-t/\tau}$ to these empirical curves, we estimate expected takedown delays $\tau\approx 31$ days for Facebook, 87 days for Instagram, 286 days for YouTube, 6 days for TikTok, and 136 days for Snapchat.
+The exponential models provide a reasonable approximation for short delay periods but diverge from the empirical curves at longer times, where heavier tails are evident.
+Consequently, these estimates of $\tau$ should be interpreted as optimistic summaries, as the exponential specification downweighs rare but prolonged takedown delays observed in the data.
+To account for both cross-platform variation and uncertainty in the tail behavior of moderation delays, our simulations (Section Takedown fails to curb illegal content prevalence and exposure) explores a broad range of $\tau$ values that both encompasses and extends beyond the empirically observed estimates, up to a maximum of $\tau=739$ days.
+Table 2: Expected content takedown delay $\tau$ of illegal content by type and platform.
+YouTube TikTok
+Illegal Content Type Observations Delay Observations Delay
+Copyright violation 375,531 276
+Privacy violation 17,737 2,362 14,174 8
+Other illegal 9,076 58 15,375 7
+Counterfeit 2,500 78
+Circumvention 1,174 769
+Hate speech 763 12 19,036 10
+Defamation 718 36 1,225 5
+Child endangerment 647 88 11,881 1
+Court order 254 174
+Non-consensual intimate imgs 12,271 6
+Fraud 4,922 54
+Harassment 4,853 4
+Terrorism 4,158 8
+National security 1,265 47
+CSAM 1,249 5
+Variation in takedown delays within platforms may further reflect differences across types of illegal content, as some violations can be assessed quickly while others require more extensive investigation.
+Table 2 reports estimated $\tau$ values by illegal content type for YouTube and TikTok.
+Corresponding empirical survival curves for each content type are shown in Appendix, Empirical expected takedown delay by illegal content type.
+The remaining three platforms are excluded from this analysis because their SoRs do not provide sufficiently informative classifications of illegal content types.
+Both YouTube and TikTok exhibit substantial heterogeneity across content categories: on YouTube, estimated delays range from 12 to 2,362 days, whereas on TikTok they range from one to 54 days.
+While this variation underscores that illegal content moderation is far from uniform, we next simulate a range of takedown scenarios designed to encompass most of the empirically observed values of $\tau$.
+Takedown fails to curb illegal content prevalence and exposure
+Based on the above empirical analysis, we simulate various takedown scenarios by exploring a broad range of expected content takedown delay values, 2 hours $\leq\tau\leq 739$ days.
+We measure the effects of time delay in content takedown by the reduction of illegal content throughout the network, as quantified by the prevalence and exposure metrics defined in Methods.
+Figure 3: Impact of takedown delay on illegal content.
+The reduction in the prevalence (pink), reach (green), and impressions (purple) of illegal content is plotted as a function of the expected takedown delay $\tau$.
+Shading represents the 95% confidence intervals estimated using non-parametric bootstrapping.
+The vertical lines indicate the expected takedown delays obtained by fitting the DSA-TDB data for different platforms.
+Fig.
+3 plots the reduction in illegal content prevalence, impressions, and reach as a function of the expected takedown delay.
+Content removal has maximal impact under rapid takedowns.
+For expected delay below nine hours, removal results in approximately 95–100% reduction of illegal content across all measures.
+Moderation becomes less effective as the expected time delay increases, allowing illegal content to spread before being removed.
+Between approximately 9 hours–46 days, the reduction in prevalence of illegal content drops from 95% to negligible levels.
+The empirical values measured from DSA-TDB data for TikTok and Facebook are within this interval.
+Those for Instagram, Snapchat, and YouTube correspond to longer delays, where removal has no significant effect on illegal content prevalence.
+Although the overall trend is similar across the two exposure metrics, the reduction in impressions and reach drops even faster for delays of up to a few days.
+For example, when illegal content is removed within a day on average, reach and impressions are only reduced by about 60%.
+In summary, these findings show that delays beyond a day significantly diminish the effect of content removal, and that all platforms have longer delays, even in our optimistic estimates.
 Discussion
-This paper reveals how delays in illegal content takedown impact its effectiveness.
+This paper reveals how delays in illegal content takedown affect prevalece and exposure of illegal content.
 Using empirical takedown data for five social media platforms from the DSA-TDB, we find large variations in average takedown delays.
-While platforms like TikTok typically respond within a week, illegal content often remains online for over a month on platforms like Facebook, Instagram, Snapchat, and YouTube.
-Our agent-based model calibrated on this data shows that long delays can substantially diminish the impact of content removal.
+While TikTok typically responds within a week, illegal content often remains online for over a month on Facebook, Instagram, Snapchat, and YouTube.
+Our agent-based model calibrated on this data shows that such long delays can substantially diminish the impact of content removal.
 Although this relationship may appear intuitive, our analysis reveals non-linear dynamics between takedown speed and moderation effectiveness, including critical delay thresholds.
 Specifically, we show that removing illegal content within one day significantly reduces its prevalence and exposure, whereas prolonged delays allow illegal content to permeate user feeds — beyond 46 days, takedown becomes futile in curbing the spread of illegal content.
-These findings have practical implications for social media moderation and regulation.
-By adopting more proactive takedown policies, platforms may reduce moderation delay and limit the spread of illegal content.
+These findings have practical implications for legislation and social media moderation.
+For legislation, the importance of timely moderation appears to support the argument for implementing tight illegal content removal deadlines, as imposed under NetzDG in Germany and KoPl-G in Austria prior to the DSA.
+However, tight takedown deadlines may have detrimental effects on moderation quality and freedom of speech.
+Takedown deadlines reduce the time necessary to properly assess the (un)lawfulness of content, which often requires careful verification, including report reviews and appeal resolutions.
+Furthermore, definitions of what content is illegal differ across countries, demanding careful assessment of national context to avoid over-removal.
+On the other hand, legislation that defines tight takedown deadlines may impose fines for platforms that systematically fail to meet them.
+The threat of such fines may incentivize platforms to over-remove content to avoid sanctions, thereby stifling freedom of speech online Europe and others (2017).
+In France, the Loi Avia was struck down by the Constitutional Council Conseil Constitutionnel (2020) precisely because of concerns about restrictions on freedom of expression Vie publique (2020); Heldt (2020).
+The Constitutional Council argued that the determination of the unlawfulness of the content is often not based on its obvious nature but is a matter for assessment by competent authorities.
+The stringent takedown deadline granted to the platforms for enforcement did not allow operators to wait for a decision by a judge, thus violating freedom of expression.
+Any regulatory framework must ensure effective moderation of illegal content while remaining practical for platforms, and mindful of potential adverse effects.
+These considerations temper the implications of our results for proactive moderation policies.
+Even though reducing takedown delays could significantly limit the spread of illegal content, platforms should remove immediately only obvious cases.
+For cases requiring further assessment (e.g., borderline content, national legislation requirements), platforms should take time for thorough review to avoid hasty decisions and over-removal.
+Until the assessment is completed, platforms may implement alternative interventions, such as downranking or reduced visibility, to reduce the circulation of potentially illegal content without resorting to premature removal.
 Another approach to improving content moderation is to strengthen collaboration with trusted flaggers — individuals or organizations that identify and report content violating community guidelines or legal standards.
-With human flaggers being a limited resource (35), efforts should be directed towards prioritizing the most time-sensitive cases.
-For legislation, the importance of timely moderation appears to support the argument for implementing tight illegal content removal deadlines, as imposed under NetzDG in Germany and KoPl-G in Austria prior to the DSA.
-However, the practical impact of such deadlines should be carefully assessed in regulatory debates for two reasons.
-First, regulatory takedown deadlines can only address the timespan between the notification of illegal content and its removal, not the period during which illegal content circulates in the network undetected.
+With human flaggers being a limited resource Facebook (2024), efforts should be directed towards prioritizing the most time-sensitive cases.
+Another issue arising from strict regulatory takedown deadlines is that they can only address the timespan between the notification of illegal content and its removal, not the period during which illegal content circulates in the network undetected.
 Therefore, the impact of tight deadlines on the spread of illegal content is inherently limited.
@@
 These figures suggest that a large portion of the long delays measured from the DSA-TDB data occurs before illegal content is detected or flagged.
-To better inform policy, platforms should be required to report not only the timestamps for publication and removal of content, but also the detection/notification timestamps to the DSA Transparency Database.
-This would allow to identify how long content has circulated in the network undetected as well as the moderation delay in response to notifications.
-Second, tight takedown deadlines may reduce the time necessary for properly assessing the (un)lawfulness of content, increasing the risk of wrong moderation decisions and thus lowering the quality of content moderation.
-Flagged content requires careful verification, including report reviews and appeal resolutions, before removal decisions can be made.
-Any regulatory framework must ensure effective moderation of illegal content while remaining practical for platforms, and mindful of potential adverse effects.
-For example, in France, the Loi Avia was struck down by the Constitutional Council (36) due to concerns about restrictions on freedom of expression (22, 23).
-These considerations suggest that platforms should remove immediately only obviously illegal content.
-For cases requiring further assessment, platforms should take time for thorough review.
-Until the assessment is completed, platforms may implement alternative interventions to reduce the circulation of potentially illegal content.
+To further increase the utility of the DSA Transparency Database for research and policy on content moderation, platforms should be required to report the timestamps not only for publication and removal of content, but also for detection/notification.
+This would make it possible to identify how long content has circulated in the network undetected, and to evaluate if platforms do react timely to violations and notifications.
+Moreover, SoRs for content and account restrictions should include additional indicators of reach and scale.
+Simple engagement metrics (number of followers, views, reshares) would enhance interpretability and support monitoring the spread of illegal content.
+Finally, individual SoRs should be linked to accounts in a privacy-preserving way to enable identification of the volume of repeat offenders Groesch et al.
+(2026).
 While our findings are robust, several limitations warrant consideration.
@@
 First, the model focuses on the spread of content already posted, without accounting for the user incentives or legal risks influencing decisions to post illegal material.
-We assume static probability distributions of sharing such content, though in reality, this behavior may evolve in response to moderation policies, potentially declining under stricter or faster moderation.
-Second, as mentioned in Methods, we assume uniform appeal and spreading potential across all types of content.
+Second, we assume static probability distributions of sharing such content.
+Future research could extend our model by accounting for the fact that users often alter their posting patterns in response to moderation decisions, potentially declining under stricter or faster moderation Srinivasan et al.
+(2019); Jhaver et al.
+(2019).
+Third, as mentioned in Methods, we assume uniform appeal and spreading potential across all types of content.
 If illegal content is generally more engaging, the model may underestimate the impact of delayed takedown, and vice versa.
-Lastly, we triangulate data from multiple sources, which may introduce inconsistencies.
-For instance, user activity is calibrated using global Twitter data, whereas the DSA database focuses on European users.
+Fourth, we triangulate data from multiple sources, which may introduce inconsistencies.
+For instance, user activity is calibrated using global Twitter data (in particular tweets linking to news), whereas the DSA-TDB focuses on European users.
 This mismatch could lead to unknown biases in our calibration.
+Lastly, our extension of SimSoMneglects out-of-network content that could be recommended in user feeds.
+Further experiments could explore the effects of out-of-network recommendations on illegal content prevalence and exposure.
 Our model of the moderation process also has limitations.
 The DSA-TDB data only captures the downstream outcomes of content moderation, neglecting the detection process.
-Our model thus treats all illegal content as equally detectable, which may overlook important variation in the detection of different content types.
+The model thus treats all illegal content as equally detectable, which may overlook important variation in the detection of different content types.
 We also make the simplifying assumption that when illegal content is removed, it is deleted from all user feeds.
 In practice, however, some content may be illegal only in certain jurisdictions and access may be selectively disabled for users in these countries.
-Our model does not consider such a differentiation.
-The database itself has limitations.
-It relies heavily on voluntary reporting by platforms, leading to inconsistent coverage and reporting quality both across and within platforms (31).
+We do not consider such a differentiation.
+The database itself has limitations Trujillo et al.
+(2025); Groesch et al.
+(2026).
+It relies heavily on self-reporting by platforms, leading to inconsistent coverage and reporting quality both across and within platforms.
 This highlights a broader issue with self-reporting systems.
-Moderation decisions are shaped by platform-specific policies and guidelines, which may not always align with legal standards: content labeled as “illegal” may in fact be lawful but removed due to internal platform rules.
-The issue of false positives may have inflated the inferred probability of illegal content from the DSA-TDB.
-Future research could extend our model by accounting for the fact that users often alter their posting patterns in response to moderation decisions (37, 38).
-One could also extend the model to compare the effects of takedown versus alternative types of interventions, as well as explore combinations of different intervention instruments.
+Moderation decisions are shaped by platform-specific moderation guidelines and practices.
+Classification of content as illegal does not follow coherent and reliable procedures across all platforms.
+Platforms are not required to assess lawfulness of content.
+Moderation is typically guided by a platform’s terms and conditions to avoid the cost and complexity of legal review Keller (2023).
+As a result, content classified as illegal may neither reflect actual unlawfulness nor capture all illegal content Groesch et al.
+(2026).
+These issues may have affected the illegal content probability inferred from the DSA-TDB.
+Future research could extend our model to compare the effects of takedown versus alternative types of interventions, as well as explore combinations of different intervention instruments.
 Finally, future models should consider the trade-offs between the harm caused by persistent illegal content and the unitended effects of moderation tools, e.g., the limitation of free speech caused by false-positive errors.
@@
 This work was supported in part by the Swiss National Science Foundation (Sinergia grant CRSII5_209250) and by the Knight Foundation.
+Appendix
+Alternative platform compliance data
+The Transparency Reports that platforms are required to publish under the DSA (Art.
+15) are a potential starting point to understand how content takedown is practiced across different platforms.
+They include metrics such as “the median time needed for taking the action.”
+However, despite formal compliance, reporting practices differ significantly across platforms, preventing thorough comparative analysis.
+Moreover, median turnaround times in transparency reports only cover the time between notification and removal of content.
+As a result, they omit the substantial period during which problematic content circulates in the network prior to notification.
+Finally, these reports do not provide moderation delays specific to illegal content, and platforms do not allow external access to the underlying data, preventing independent validation.
+To overcome these limitations, we instead rely on the DSA Transparency Database (DSA-TDB), as described in the main text.
+Illegal content probability modeling
+As there is limited access to data on illegal content, it is difficult to determine how such individual probability of posting illegal content is distributed across users.
+We consider three alternative distributions of individual-level illegal posting probabilities that all share the same population mean $p=0.01$.
+First, we consider a nearly homogeneous scenario in which all users have similar propensities to post illegal content.
+In this case, $p$ is drawn from a normal distribution with small variance, $P(i)\sim\mathcal{N}(\mu=0.01,\sigma=0.001)$.
+Second, we consider a skewed but unimodal distribution that allows for moderate heterogeneity.
+We model this using a beta distribution, $P(i)\sim i^{\alpha-1}(1-i)^{\beta-1}$, where the mean is given by $\frac{\alpha}{\alpha+\beta}$.
+We set $\alpha=10$ and $\beta=990$, yielding the same mean probability $p=0.01$.
+Third, we explicitly model population heterogeneity by distinguishing between high-risk and low-risk accounts.
+This reflects the intuition that a small fraction of users may repeatedly post severe illegal content (e.g., CSAM, terrorist propaganda, or scams), while the majority of users only occasionally violate the law, for example by sharing copyrighted material.
+In this two-group model, a fraction $s_{H}$ of users belong to the high-risk group, while the remaining $1-s_{H}$ belong to the low-risk group.
+Within each group, individual probabilities are again drawn from beta distributions as described above.
+The average illegal content probability across the full population is thus $p=s_{H}\frac{\alpha_{H}}{\alpha_{H}+\beta_{H}}+(1-s_{H})\frac{\alpha_{L}}{\alpha_{L}+\beta_{L}}$, where the $\alpha$ and $\beta$ subscripts indicate the groups.
+We set $s_{H}=0.1$ to reflect a minority of high-risk accounts.
+We use the values of $\alpha_{L}=0.1$, $\beta_{L}=90$ for the low-risk group and $\alpha_{H}=3$, $\beta_{H}=30$ for the high-risk group, resulting in $p=0.01$.
+We explore whether the main outcome variable is sensitive to these different assumptions about illegal content distributions.
+The reduction in illegal content prevalence shows no statistically significant differences, suggesting that the choice of underlying distribution does not affect the main findings (see Fig.
+S1 in Robustness).
+The simulations discussed in the main text use the two beta distributions.
+We additionally show that our findings are robust to the relative sizes of the groups (see Fig.
+S2 in Robustness).
+Figure S1: Robustness of results with respect to different distributions of $p$ for $\tau=2.89$ (median takedown delay of two days, top) and $\tau=11.54$ (median takedown delay of eight days, bottom) based on 10 simulations.
+The illegal content reduction does not vary significantly ($P\geq 0.13$ across pairs of simulations with different distributions, using Mann–Whitney U tests with Bonferroni correction).
+Robustness
+The results presented in the main text may depend on the distributions of individual-level illegal posting probabilities, the proportion of high-risk accounts $s_{H}$, the illegal content probability $p$, and the underlying follower network.
+We therefore investigate if the results remain valid when changing these parameters.
+For these robustness checks, we use two different values of expected content takedown delay, $\tau=2.89$ and $\tau=11.54$, corresponding to median delays of 2 and 8 days, respectively, and assign users to the high- and low-risk groups randomly in each run.
+Fig. S1 shows that the main results are robust with respect to different assumptions about the distributions of individual illegal posting probabilities.
+For the following analyses varying high-risk group size and illegal content probability, we need a large follower network to generate stable interaction dynamics using low values of these parameters ($s_{H}=10^{-3}$ and $p=3\times 10^{-4}$).
+Therefore, we use the original network but remove a random subset of edges to speed up the simulations.
+The average in/out-degree is thus reduced to 18 friends per node, resulting in a network with N=58,296N=58{,}296 nodes and E=1,063,202E=1{,}063{,}202 edges.
+Figure S2: Robustness of results with respect to different high-risk group sizes $s_{H}$ for $\tau=2.89$ (median takedown delay of two days, top) and $\tau=11.54$ (median takedown delay of eight days, bottom) based on 20 simulations.
+The illegal content reduction does not vary significantly ($P\geq 0.72$ across pairs of simulations with different $s_{H}$ values, using Mann–Whitney U tests with Bonferroni correction).
+Let us examine the sensitivity of the results to the size of the high-risk group.
+While our baseline setting assumes $s_{H}=0.1$, we explore several orders of magnitude smaller values of the high-risk group size ($10^{-3}\leq s_{H}\leq 10^{-1}$) to capture scenarios in which illegal content production is concentrated among a very small fraction of users.
+The main outcome variable, reduction in illegal content prevalence, shows no statistically significant differences (Fig.
+S2), suggesting that our findings are robust to changes in $s_{H}$.
+We wish to explore how the main result is affected by the volume of illegal content, which varies across platforms and countries.
+We run simulations with the empirical estimates of $p$ reported in Table 1 (except for the smallest value, which would require a network with infeasibly large size).
+In the main text, we report the reduction in illegal content prevalence, $1-\frac{\bar{I}_{\text{removal}}}{\langle\bar{I}_{\text{baseline}}\rangle}$.
+To show that this ratio is independent of $p$, we plot the scaling relationship between $\bar{I}_{\text{removal}}$ and $\bar{I}_{\text{baseline}}$ in Fig.
+S3a for different $p$ values from Table 1.
+We observe only a weak sublinear relationship, $\bar{I}_{\text{removal}}\sim\bar{I}_{\text{baseline}}^{0.94}$, suggesting that the ratio is nearly independent of $p$.
+We also plot in Fig.
+S3b the reduction in the prevalence of illegal content for different values of $p$ from Table 1.
+We observe more noise for small $p$, but none of the differences between the values in our main results ($p=0.01$) and those for smaller $p$ values are statistically significant.
+These analyses suggest that our main results are robust with respect to the volume of illegal content in the system.
+Figure S3: Robustness analysis for illegal content probability.
+(a) Scaling between illegal content prevalence in removal and baseline conditions.
+Each data point represents the proportion of illegal content in the system with and without removal.
+Results are based on 40 simulations using $\tau=2.89$ (median takedown delay of two days).
+Each simulation is done with an empirical estimate of $p$ (vertical lines) for Facebook and YouTube users from Germany, France, and Sweden, derived from the data reported in Table 1.
+Estimates of $p$ are based on the proportion of disappeared content that is subsequently judged to be legal.
+The gray line represents a least-squares fit of the log-transformed data, yielding a slope of $0.94\pm 0.01$.
+(b) Illegal content reduction for different illegal content probabilities corresponding to the same countries and platforms, based on 10 simulations with $\tau=2.89$ (median takedown delay of two days, top) and $\tau=11.54$ (median takedown delay of eight days, bottom).
+The result does not vary significantly ($P\geq 0.13$ across pairs of simulations with different $p$ values, using Mann–Whitney U tests with Bonferroni correction).
+Last, we examine the effect of network structure by comparing simulations using the empirical network and a synthetic network.
+We create a synthetic network with features ubiquitous in real-world networks, namely, the presence of hubs and clustering (directed triads).
+To this end, we use a directed variant of the random-walk growth model Vázquez (2003).
+We configure the parameters of this model to generate a network comparable in size and average degree to the network used in the main experiment.
+In particular, we initialize the network with 181 fully connected nodes and assume new nodes have fixed out-degree $k_{out}=180$.
+New nodes are added by first randomly selecting one node to follow.
+For the remaining $k_{out}-1$ connections, each new node has an equal probability to follow a random friend of the initially chosen node or a randomly selected node in the network.
+This results in a network with N=10,006N=10{,}006 nodes and E=1,810,905E=1{,}810{,}905 edges.
+There is no statistically significant difference in the reduction in illegal content prevalence between these two networks, indicating the robustness of our findings (Fig.
+S4).
+Figure S4: Robustness analysis with respect to network structures based on 10 simulations with $\tau=2.89$ (median takedown delay of two days, left) and $\tau=11.54$ (median takedown delay of eight days, right).
+The illegal content reduction does not vary significantly ($P=0.11$ and $P=0.22$, respectively, using Mann–Whitney U tests with Bonferroni correction).
+Empirical expected takedown delay by illegal content type
+Figure S5: Distributions (CCDF) of time delay in takedown on YouTube by type of illegal content.
+Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line).
+Figure S6: Distributions (CCDF) of time delay in takedown on TikTok by type of illegal content.
+Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line).
+We wish to examine the variation in takedown delays across types of illegal content (Table 2).
+Fig.
+S5 and Fig.
+S6 show empirical survival time curves for different illegal content types from DSA-TDB SoR records of YouTube and TikTok.
+By fitting exponential survival time curves $P(t)\sim e^{-t/\tau}$ to this empirical data, we estimate the ranges of $\tau$ for these platforms: 12 days $\leq\tau\leq$ 2,362 days for YouTube and 1 day $\leq\tau\leq$ 54 days for TikTok.